### v1

# AdaGC: Improving Training Stability for Large Language Model Pretraining

Abstract
Large Language Models (LLMs) face increasing loss spikes during scaling, undermining training stability and final performance. While gradient clipping mitigates this issue, traditional global approaches poorly handle parameter-specific gradient variations and decaying gradient norms. We propose AdaGC, an adaptive gradient clipping framework that automatically adjusts local thresholds per parameter through exponential moving average of gradient norms. Theoretical analysis proves AdaGC’s convergence under non-convex conditions. Extensive experiments demonstrate significant improvements: On Llama-2 7B/13B, AdaGC completely eliminates loss spikes while reducing WikiText perplexity by 3.5% (+0.14pp LAMBADA accuracy) for 7B and achieving 0.65% lower training loss with 1.47% reduced validation perplexity for 13B compared to global clipping. For CLIP ViT-Base, AdaGC converges 25% faster than StableAdamW with full spike elimination. The method shows universal effectiveness across architectures (Llama-2 7B/13B) and modalities (CLIP), with successful integration into diverse optimizers like AdamW and Lion. Source code will be released on GitHub.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b) have become a cornerstone of modern Natural Language Processing (NLP) research, demonstrating outstanding performance in a wide range of language tasks. These models, characterized by their extensive parameter counts and vast training datasets, are designed to capture the subtleties of human language. However, as these models grow in size and complexity, they often encounter a significant training obstacle: the loss spike—a sharp, unexpected increase in training loss that can seriously undermine the training process’s efficiency and effectiveness (Chowdhery et al., 2023; Scao et al., 2022; Takase et al., 2023; Zeng et al., 2022).
Among the various methods for addressing spike issues, gradient clipping is an effective solution. Existing gradient clipping approaches (Pascanu et al., 2013) typically apply a predefined threshold to the gradient norm of the entire neural network. However, we have identified two significant shortcomings with this global threshold method. First, as the gradient magnitudes decrease during training, an excessively small threshold throughout the process can slow down the training speed, while a consistently large threshold fails to effectively mitigate spike issues in the later stages of training. Second, due to the inherent differences between structures in neural networks (Zhang et al., 2024), particularly in transformers, applying a uniform clipping coefficient across all layers does not adequately resolve the problems caused by spikes, as illustrated in the Figure 1. To address these two issues with global gradient clipping, we propose using local gradient clipping along with an adaptive update of the clipping threshold for each parameter.
Specifically, we introduce a novel method called Adaptive Gradient Clipping based on Local Gradient Norm (AdaGC). AdaGC is specifically designed to mitigate loss spike by selectively suppressing abnormal gradients, thereby maintaining training stability and enhancing model performance. At its core, AdaGC employs the Exponential Moving Average (EMA) to accurately track historical gradient norm variations at the tensor level. This enables the precise identification and adaptive clipping of anomalous local gradients, effectively containing the loss spike.
Convergence analysis shows AdaGC maintains the $\mathcal{O}(1/\sqrt{T})$ rate comparable to Adam in non-convex settings, confirming that localized clipping preserves convergence properties. Comprehensive experiments on Llama-2 7B/13B (Touvron et al., 2023b) and CLIP ViT-Base (Radford et al., 2021) demonstrate complete loss spike elimination while improving model convergence (3.5% perplexity reduction, 25% faster convergence than StableAdamW (Wortsman et al., 2023a)), with successful integration into modern optimizers like AdamW (Loshchilov & Hutter, 2017) and Lion (Chen et al., 2024).
In summary, our main contributions are three-fold:
  • Identification of two critical limitations in global gradient clipping methods: inadequate handling of decaying gradient norms and parameter-wise heterogeneity, motivating the need for localized adaptive clipping.
  • Proposal of Adaptive Gradient Clipping (AdaGC), featuring parameter-wise threshold adaptation through exponential moving averages, effectively addressing both temporal gradient decay and spatial parameter variations.
  • Comprehensive empirical validation across architectures (Llama-2 7B/13B) and modalities (CLIP ViT-Base), demonstrating complete loss spike elimination with measurable performance gains, plus convergence analysis confirming comparable rates to standard optimizers.
2 Related Work
Loss spikes present a significant challenge when scaling models, potentially impeding learning or destabilizing the training process. Various strategies have been proposed to enhance the stability of pre-training in large language models.
Training Strategies: Effective training strategies are pivotal for achieving stability. Adopting the bfloat16 format (Wang & Kanwar, 2019) is a cornerstone for stable training. Using smaller learning rates (Wortsman et al., 2023b; Zhang et al., 2022) also promotes stable training dynamics. To mitigate loss spikes, PaLM (Chowdhery et al., 2023) resumes training from a checkpoint approximately 100 steps before the spike and bypasses 200-500 data batches to avoid those preceding and during the spike. Varying the sequence length (Li et al., 2022) is also effective for stabilizing the pre-training of large language models.
Model Architecture: Advanced architectures are crucial for stability in training large-scale models. Research shows that Pre-LN Transformers offer enhanced stability over Post-LN Transformers (Xiong et al., 2020; Vaswani et al., 2017; Liu et al., 2020), both theoretically and practically. Thus, contemporary studies predominantly employ Pre-LN Transformers for building large-scale models. Techniques such as NormHead (Yang et al., 2023), which normalizes output embeddings, and RMSNorm (Zhang & Sennrich, 2019), as reported in Llama (Touvron et al., 2023a), contribute to stable training. Applying layer normalization to the embedding layer (Scao et al., 2022), normalizing the input vector before the standard softmax function (NormSoftmax) (Jiang et al., 2023), and the shrink embedding gradient technique (Zeng et al., 2022) are also beneficial.
Model Initialization: Proper model initialization significantly enhances training stability. Scaling the embedding layer with an appropriate constant (Takase et al., 2023) reinforces LLM pre-training stability. Initializing Transformers with small parameter values (Nguyen & Salazar, 2019) also improves pre-training stability. Some researchers propose that removing layer normalizations in Transformers is feasible with specific initialization methods (Zhang et al., 2019; Huang et al., 2020).
Auxiliary Loss Function: Incorporating an auxiliary loss function alongside conventional language modeling loss enhances training stability. Notable examples include Max-z loss, which benefits the pre-training process (Chowdhery et al., 2023; Wortsman et al., 2023b; Yang et al., 2023).
Gradient/Update Clipping: Gradient and update clipping achieve stability by limiting the magnitude of parameter updates, preventing excessively large weight updates. Global gradient clipping (Pascanu et al., 2013) is prevalent, with innovative approaches like adaptive gradient clipping (Brock et al., 2021) and Clippy (Tang et al., 2023), which use model weights to adjust the clipping threshold. Alternatives like Adafactor (Shazeer & Stern, 2018), StableAdamW (Wortsman et al., 2023a), and LAMB (You et al., 2019) offer update clipping techniques better suited for stability training of large-scale models. Nonetheless, a significant number of loss spikes still occur during the training of large language models, even with the application of these methodologies.
3 Motivations
(a) Temporal threshold adaptation necessity: Fixed thresholds (0.4) prevent early spikes but fail later (10k steps), requiring dynamic adjustment.
(b) Spikes in gradient norms occur at different training steps across parameters, indicating heterogeneous gradient behaviors.
(c) Per-parameter gradient clipping mitigates loss spikes that global clipping cannot, leading to faster convergence.
Figure 1: Empirical motivation for AdaGC: (a) Temporal threshold decay necessitates adaptive clipping, (b) Parameter-specific gradient spikes demand localized control, (c) Fine-grained clipping outperforms global approaches.
In this section, we will introduce some phenomena of gradient clipping that motivate the design of the algorithm.
Adaptivity: The magnitudes of gradients in neural network training exhibit dynamic variations throughout the learning process. This renders fixed gradient clipping thresholds suboptimal for mitigating loss spikes, as demonstrated in Figure 1(a). While a threshold of 0.4 successfully mitigates loss spike during the initial training phase of the GPT-2 345M model, it becomes ineffective against the severe loss spike occurring around 10,000 iterations, where a substantially reduced threshold is required for stabilization. This phenomenon correlates with the observation that optimal clipping thresholds for convergence speed are strongly influenced by recent gradient norms (see corresponding subfigures). Motivated by these findings, we propose an adaptive clipping mechanism using the exponential moving average of gradient magnitudes as the per-step clipping threshold.
Locality: Figure 1(b) reveals that different parameters exhibit divergent gradient behaviors, with spikes occurring at distinct training phases. Using a global clipping threshold can degrade the convergence speed of certain parameters. Therefore, we propose replacing global gradient norm clipping with per-parameter norm clipping, where clipping is applied independently to each parameter. Furthermore, as illustrated in Figure 1(c), we evaluate the effectiveness of per-parameter clipping thresholds categorized by specific ratios on GPT2-345M. The results indicate that, under the same global gradient norm conditions, per-parameter clipping can address spike phenomena that global clipping cannot resolve, and achieve faster convergence speeds.
These observations collectively motivate our two core design principles: adaptive thresholding to address temporal gradient norm decay, and localized control to handle parameter-wise gradient heterogeneity. The dynamic nature of optimal clipping thresholds (Figure 1(a)) necessitates our exponential moving average mechanism for temporal adaptation, while asynchronous parameter gradient spikes (Figure 1(b)) justify per-parameter clipping granularity. Together, these principles form the foundation of AdaGC - adaptively adjusting localized thresholds that respect both the temporal evolution and spatial distribution of gradients.
4 Methodology: AdaGC
4.1 Preliminaries
[…]
Gradient Clipping Fundamentals. Consider a stochastic optimization problem with parameters $\theta\in\mathbb{R}^{d}$ and loss function $f(\theta;X_{t})$ evaluated on mini-batch $X_{t}$ at step $t$. Standard gradient descent updates follow:
$\theta_{t}=\theta_{t-1}-\alpha\nabla_{\theta}f(\theta_{t-1},X_{t})$ (1)
To prevent unstable updates from gradient explosions, gradient clipping (Pascanu et al., 2013) modifies the update rule as:
$\begin{array}[]{c}\theta_{t}=\theta_{t-1}-\alpha h_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})\\[3.99994pt] \text{ where }h_{t}:=\min\left\{\frac{\lambda_{abs}}{\|\nabla_{\theta}f(\theta_{t-1};X_{t})\|},1.0\right\}\end{array}$ (2)
Here $\lambda_{\text{abs}}$ is an absolute clipping threshold requiring careful tuning. Our work focuses on norm-based clipping (scaling entire gradients exceeding $\lambda_{\text{abs}}$) rather than value-based clipping (element-wise truncation).
4.2 Adaptive Gradient Clipping based on Local Gradient Norm
This section introduces a novel gradient clipping strategy termed Adaptive Gradient Clipping (AdaGC), which distinguishes itself by not relying on a global gradient norm. Instead, AdaGC focuses on the local gradient norm of each parameter and utilizes a dynamic adaptive mechanism for gradient clipping. The proposed method employs an Exponential Moving Average (EMA) mechanism to maintain smoothed estimates of historical gradient norms per parameter, thus enhancing the accuracy of anomalous gradient detection and enabling independent clipping adjustments tailored to each parameter’s specific conditions. EMA is widely used in deep learning, and within AdaGC, it facilitates the balancing of historical and current gradient norms. The formulation is as follows:
$\begin{array}[]{c}\boldsymbol{g}_{t,i}\leftarrow h_{t,i}\cdot\boldsymbol{g}_{t,i},\text{where }h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\},\\[10.00002pt] \gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|.\end{array}$ (3)
Here, $\lambda_{rel}$ is a predefined relative clipping threshold, $\boldsymbol{g}_{t,i}$ represents the gradient of the $i$-th parameter at time step $t$, and $h_{t,i}$ is a clipping function activated when $\|\boldsymbol{g}_{t,i}\|>\lambda_{rel}\cdot\gamma_{t-1,i}$, thereby scaling the gradient norm to $\lambda_{rel}\cdot\gamma_{t-1,i}$. Additionally, $\beta$ is the smoothing coefficient for EMA. We consistently incorporate the clipped gradient norm into the historical observations rather than the pre-clipped values.
Despite its simplicity, AdaGC adaptively adjusts based on the magnitude of each parameter’s gradient norm. Whenever the gradient norm at a current timestep exceeds a predefined range of average norms within a historical window, it effectively suppresses these outlier gradients.
However, during the initial stages of model training (e.g., the first 100 steps), the gradient norms are typically large and fluctuate significantly, indicating a substantial decreasing trend. Direct application of AdaGC during this period could lead to two issues: first, erroneously accumulating the early large gradient norms into the historical values, resulting in compounded errors; second, compared to traditional global-norm-based methods, AdaGC might delay clipping, thus potentially slowing down the loss reduction. To address these issues, we introduce a hyperparameter $T_{start}$(default set to 100), representing a warm-up period during which traditional global gradient norm clipping is applied.
Additionally, the AdaGC strategy can be seamlessly integrated with various optimizers, such as AdamW (Loshchilov & Hutter, 2017), enhancing its practicality and flexibility. Algorithm 1 demonstrates its implementation with the AdamW optimizer.
Overall, AdaGC offers a more precise and effective gradient management method for large-scale model training, contributing to improved training stability and performance.
Algorithm 1 AdamW with AdaGC
1:  given: $\{\alpha_{t}\}_{t=1}^{T},\lambda_{w},\epsilon_{1},\beta_{1},\beta_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}\in[0,1),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{abs}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{start}}$
2:  initialize: $\boldsymbol{\theta}_{0},m_{0}\leftarrow 0,v_{0}\leftarrow 0,t\leftarrow 0$
3:  repeat
4:    compute $\boldsymbol{g}_{t}$ = $\nabla_{\boldsymbol{\theta}}f_{t}(\boldsymbol{\theta}_{t-1},X_{t})$
5:    if $t<T_{start}$ then
6:       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h_{t}=\min\left\{\frac{\lambda_{abs}}{\|\boldsymbol{g}_{t}\|},1.0\right\}}$
7:       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}=h_{t}\cdot\boldsymbol{g}_{t}}$
8:       for $i\in|\boldsymbol{\theta}|$ do
9:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\min\left\{\gamma_{t-1,i},\|\widehat{\boldsymbol{g}_{t,i}}\|\right\},\gamma_{0,i}=\|\widehat{\boldsymbol{g}_{1,i}}\|}$
10:       end for
11:    else
12:       for $i\in|\boldsymbol{\theta}|$ do
13:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\}}$
14:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t,i}}=h_{t,i}\cdot\boldsymbol{g}_{t,i}}$
15:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\widehat{\boldsymbol{g}_{t,i}}\|}$
16:       end for
17:    end if
18:    $\boldsymbol{m}_{t}=\beta_{1}\boldsymbol{m}_{t-1}+(1-\beta_{1}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}}$
19:    $\boldsymbol{v}_{t}=\beta_{2}\boldsymbol{v}_{t-1}+(1-\beta_{2}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}}^{2}$
20:    $\widehat{\boldsymbol{m}_{t}}=\boldsymbol{m}_{t}/(1-\beta_{1}^{t}),\quad\widehat{\boldsymbol{v}_{t}}=\boldsymbol{v}_{t}/(1-\beta_{2}^{t})$
21:    $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\alpha_{t}\lambda_{w}\boldsymbol{\theta}_{t-1}-\alpha_{t}\widehat{\boldsymbol{m}_{t}}/(\sqrt{\widehat{\boldsymbol{v}_{t}}}+\epsilon_{1})$
22:  until $\boldsymbol{\theta}_{t}$ not converge
4.3 Convergence Analysis
In this section, we give the convergence guarantee of Adam with AdaGC stated as follows:
Theorem 4.1.
Under mild assumptions, by selecting $\alpha_{t}=\mathcal{O}(1/\sqrt{T})$, $\beta_{2}=1-\mathcal{O}(1/T)$ and $\beta_{1}<\sqrt{\beta_{2}}$, when $\tau$ is randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities, it holds that
$\mathbb{E}\|\nabla f(\theta_{\tau})\|^{2}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
Theorem 4.1 shows that even with local clipped gradient, Adam with AdaGC can converge in the same rate as vanilla Adam. Due to the limited space the formal assumptions, theorem statement with detailed proof can be found in Appendix A.
5 Experiments
This paper primarily focuses on the pre-training of large language models, using Llama-2 (Touvron et al., 2023b) as our primary experimental subject. Experiments were conducted using NVIDIA GPU A100 40G. Moreover, our proposed method demonstrates strong generalization across various VLM models, achieving robust results.
Table 1: Zero-shot performance of Llama-2 7B on WikiText and LAMBADA datasets.
Metrics GlobalGC AdaptiveGradientClip StableAdamW AdaGC-Shard AdaGC
WikiText Word PPL $\downarrow$ 21.173 21.467 21.733 20.878 20.434
Byte PPL $\downarrow$ 1.770 1.774 1.778 1.765 1.758
LAMBADA ACC $\uparrow$ 49.35 ($\pm$ 0.70) 49.27 ($\pm$ 0.70) 48.96 ($\pm$ 0.70) 47.76 ($\pm$ 0.70) 49.49 ($\pm$ 0.70)
PPL $\downarrow$ 11.111 ($\pm$ 0.334) 11.509 ($\pm$ 0.354) 11.774 ($\pm$ 0.363) 11.518 ($\pm$ 0.341) 10.515 ($\pm$ 0.315)
5.1 Experimental Setup
Models and Datasets: We evaluated the efficacy of AdaGC on various Llama-2 models including Tiny, 7B and 13B. The C4-en (Raffel et al., 2020) dataset, a clean corpus of English text extracted from Common Crawl, served as our pre-training dataset. We assessed model performance by computing perplexity on the WikiText (Merity et al., 2016) and LAMBADA (Paperno et al., 2016) datasets. We also conducted experiments on GPT-2 (Radford et al., 2019) and CLIP (Radford et al., 2021) models to further validate AdaGC’s generalization capabilities.
Comparison Methods: We selected several techniques including Global Gradient Clipping (GlobalGC for short) (Pascanu et al., 2013), Adaptive Gradient Clipping (Brock et al., 2021), StableAdamW (Wortsman et al., 2023a), and Scaled Embed (Takase et al., 2023) from different approaches such as gradient clipping, update clipping, and parameter initialization for comparison.
Training Details: Pre-training large scale models, particularly Llama-2 7B and 13B, is typically resource-intensive. Like StableAdamW (Wortsman et al., 2023a) and Scaled Embed (Takase et al., 2023), our primary focus was to explore training instability rather than achieve ultimate accuracy. For ease of multiple experiments, we conducted 9,000 training steps on 36 billion tokens for both Llama-2 7B and 13B models, and 36,000 steps on 36 billion tokens for the Llama-2 Tiny model. For additional details on the hyperparameters, please refer to the Appendix B.
Critical Hyperparameter Selection. We conducted a systematic study to evaluate the two critical hyperparameters in AdaGC: the EMA coefficient $\beta$ and the relative clipping threshold $\lambda_{rel}$. Our analysis utilized coordinate descent (Bertsekas, 1997) for hyperparameter optimization on the Llama-2 Tiny model. Initially, we examined the relative clipping threshold $\lambda_{rel}$ with a fixed EMA coefficient $\beta=0.98$, testing the values $\{1.01,1.05,1.10\}$. As depicted in Figure 2(a), $\lambda_{rel}=1.05$ achieved optimal performance by effectively balancing gradient regulation effectiveness (compromised at $\lambda_{rel}\geq 1.10$) and update stability (degraded at $\lambda_{rel}\leq 1.01$). With this optimal threshold established, we proceeded to investigate the EMA smoothing factor $\beta\in\{0.95,0.98,0.99\}$. Figure 2(b) demonstrates that $\beta=0.98$ provides superior adaptation speed while maintaining historical consistency, yielding the best performance. Consequently, we adopted the parameter settings $\lambda_{rel}=1.05$ and $\beta=0.98$ for all subsequent experiments reported in this paper.
(a) Relative clipping threshold $\lambda_{\text{rel}}$.
(b) EMA coefficient $\beta$.
Figure 2: Study of AdaGC’s hyperparameters on Llama-2 Tiny. (a) Relative clipping threshold analysis demonstrates $\lambda_{\text{rel}}=1.05$ achieves optimal gradient regulation. (b) EMA coefficient analysis reveals $\beta=0.98$ best balances historical consistency with rapid adaptation.
5.2 Experimental Results
(a) Llama-2 7B training dynamics: AdaGC (green) eliminates loss spikes observed in baseline methods (GlobalGC-blue, ScaledEmbed-orange, AdaptiveGradientClip-purple, StableAdamW-cyan), achieving final training loss 0.02 lower than GlobalGC. Validation perplexity (right) shows correlated improvement with stable gradient norms (middle).
(b) Llama-2 13B scalability analysis: AdaGC maintains superior stability at scale, reducing training loss and validation perplexity by 0.64% and 1.47% respectively versus GlobalGC, demonstrating method scalability.
Figure 3: Large language model training analysis: (a) 7B model comparison shows AdaGC’s loss spike elimination and performance gains, (b) 13B results demonstrate method scalability.
Our comprehensive evaluation demonstrates AdaGC’s effectiveness across model scales and architectures. Figure 3 presents comparative results on Llama-2 7B and 13B models, analyzing training dynamics through loss trajectories, gradient norms, and validation perplexities. For the 7B model, baseline methods (GlobalGC, Scaled Embed, Adaptive Gradient Clipping, and StableAdamW) exhibit frequent loss spikes during training, while AdaGC completely eliminates these instability events. This stability translates to measurable performance gains, with AdaGC achieving a 0.02 lower final training loss (2.28 vs 2.30) and 2.1% reduced validation perplexity compared to the strongest baseline (GlobalGC).
The approach demonstrates strong scalability on the Llama-2 13B model, where AdaGC eliminates loss spikes entirely compared to GlobalGC. Qualitative analysis reveals AdaGC’s superior outlier gradient management, maintaining smoother loss trajectories throughout training. This enhanced stability yields concrete performance improvements: 0.65% lower training loss (absolute difference of 0.0146) and 1.47% reduced validation perplexity relative to GlobalGC.
Downstream zero-shot evaluation on WikiText and LAMBADA datasets (Table 1) confirms the practical benefits of stable training. AdaGC achieves state-of-the-art performance across all benchmarks, reducing WikiText perplexity by 3.5% (20.434 vs 21.173) while improving LAMBADA accuracy by 0.14 percentage points (49.49% vs 49.35%) compared to GlobalGC. These results establish a direct correlation between training stability and final model quality.
To validate architectural generality, we extend evaluation to vision-language pretraining with CLIP ViT-Base. As shown in Figure 4(a), AdaGC completely eliminates loss spikes observed in vision-specific stabilization methods while achieving 25% faster convergence than StableAdamW (reaching target loss in 15k vs 20k steps). The improved training dynamics translate to superior zero-shot recognition performance (Figure 4(b)), with AdaGC demonstrating consistent accuracy gains across modalities - including a 0.27 percentage point improvement on ImageNet-1K (39.84% vs 39.57%) over GlobalGC.
Figure 4: CLIP ViT-Base results: (a) Training loss trajectory comparison. AdaGC achieves 25% faster convergence (15k vs 20k steps) with complete loss spike elimination compared to StableAdamW. (b) Zero-shot recognition performance. Final accuracy improves 0.27pp (39.84% vs 39.57%) on ImageNet-1K, demonstrating cross-modal generalization.
5.3 Ablation Study
We conduct systematic ablation studies across three critical dimensions of AdaGC: (1) EMA gradient norm initialization strategies, (2) adaptivity efficacy, and (3) locality granularity. The adaptivity property has been empirically validated through experimental analysis (Section 3) and comprehensive experimental results across all benchmarks.
EMA Initialization Strategy.
The initialization of EMA gradient norms requires careful design due to large initial gradient fluctuations during early training phases (first 100 steps). We evaluate five initialization variants on GPT-2 345M: The default AdaGC strategy employs GlobalGC during warm-up while tracking minimum per-parameter norms ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$). Comparative approaches include: (1) norm initialization without GlobalGC warm-up (directly using $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$ from step 0), (2) constant initialization ($\gamma_{0,i}\in\{0.5,1.0\}$), and (3) thresholded initialization ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,0.1)$). Figure 6(b) demonstrates that while all variants eliminate loss spikes, convergence quality varies within 0.36%. The default strategy achieves optimal final loss (2.9708 vs 2.9725 for next-best), showing that GlobalGC-guided warm-up better preserves parameter update consistency than direct initialization. This establishes the importance of phased initialization for gradient norm adaptation.
Locality Granularity. We analyze clipping granularity across three implementation levels. Global-level clipping applies a single threshold for the entire model, while parameter-wise adaptation operates per individual parameter. On GPT-2 345M (Figure 5(a)), global clipping reduces but fails to eliminate spikes (1 event vs 0 for parameter-wise), with 0.25% (2.9639 vs 2.9712) higher final loss.
As model scale increases to 7B parameters, TensorParallel distributed training becomes necessary for efficient optimization. We therefore evaluate shard-wise adaptation on Llama-2 7B, where parameters are partitioned across devices and clipping operates per-shard. Figure 5(b) reveals that while shard-wise clipping eliminates spikes, it introduces parameter update inconsistencies across devices, increasing loss variance by 4.3% (2.3836 vs 2.28) compared to parameter-wise granularity. This degradation stems from independent clipping of parameter shards disrupting holistic parameter updates.
These results establish parameter-wise granularity as optimal, achieving precise adaptation while maintaining parameter update coherence across distributed systems.
Figure 5: Locality granularity analysis: (a) Global vs parameter-wise clipping on GPT-2 345M model demonstrates spike elimination capability, (b) Distributed shard-wise clipping on 7B model reveals parameter integrity requirements.
Figure 6: (a) Large rate tolerance on GPT-2 345M: AdaGC enables stable training at 10$\times$ learning rate (left), outperforms baselines (mid), and accelerates convergence through rate scaling (right). (b) EMA initialization strategies.
5.4 Exploring AdaGC’s Capabilities
We demonstrate AdaGC’s potential advantages over global gradient clipping through three critical dimensions: large learning rate tolerance, batch size scalability, and optimizer compatibility.
Large Learning Rate Tolerance
Figure 6(a) analyzes training stability under aggressive learning rate regimes using GPT-2 345M. At $3\times 10^{-3}$ learning rate (10$\times$ baseline), AdaGC maintains stable training (0 loss spikes) while achieving 2.52% lower final loss than GlobalGC (left panel). Comparative analysis with stabilization baselines (middle panel) reveals AdaGC’s unique capability: 1.52% lower training loss than GlobalGC+ScaledEmbed (best competitor) without spike mitigation overhead. The scaling law analysis (right panel) suggests a noticeable convergence acceleration with increasing learning rates (from $3\times 10^{-4}$ to $3\times 10^{-3}$). This indicates that learning rate scaling may contribute to a faster training process, albeit not with a perfect linear relationship.
Batch Size Scalability
Figure 7 validates AdaGC’s effectiveness under scaled batch training paradigms. With batch sizes up to 8192 (16$\times$ baseline) and proportional learning rate scaling, AdaGC maintains perfect stability (0 spikes vs 100% spike rate for GlobalGC) while reducing final perplexity by 1.28% - 19.55% across configurations. The Llama-2 Tiny experiments confirm cross-architecture effectiveness, showing 6.19% lower perplexity than GlobalGC at a batch size of 8192. This demonstrates AdaGC’s capability to enable efficient large-batch distributed training without stability compromises. Additional details can be found in Figure 9 and Figure 10 in Appendix D.
(a) bs = $512$, lr = $3.0\times 10^{-4}$
(b) bs = $2048$, lr = $6.0\times 10^{-4}$
(c) bs = $4096$, lr = $8.5\times 10^{-4}$
(d) bs = $8192$, lr = $1.2\times 10^{-3}$
Figure 7: AdaGC’s large-batch scalability on GPT-2 345M: Maintains 0 loss spikes (vs GlobalGC’s 100% rate) with batch sizes up to 8192 (16$\times$ baseline).
Optimizer Compatibility
Figure 8 demonstrates AdaGC’s generalization across optimization frameworks. Integrated with Lion, AdaGC achieves comparable training loss to GlobalGC (1.9912 vs 2.0165) while improving ImageNet-1K top-1 accuracy by 0.16 percentage points (40.81% vs 40.65%). When combined with StableAdamW (which explicitly avoids gradient clipping), AdaGC provides additional 19.23% training loss reduction (2.0523 vs 2.5412) and 0.96 percentage point accuracy gain (39.71% vs 38.75%), establishing complementary benefits. These results position AdaGC as a universal stabilization component for modern optimization paradigms.
Figure 8: AdaGC’s optimizer compatibility on CLIP ViT-Base: With Lion, achieves 0.16pp higher accuracy (40.81% vs 40.65%) at comparable loss; Combined with StableAdamW, reduces loss by 19.23% (2.0523 vs 2.5412) with 0.96pp accuracy gain. Demonstrates universal stabilization across optimization frameworks.
6 Discussion
6.1 Socio-Economic Value
Pre-training large models entails substantial costs and is vulnerable to significant economic losses due to instabilities in the training process. Loss spikes, typically addressed by resuming training from the latest checkpoint as illustrated by the PaLM (Chowdhery et al., 2023) approach, can have severe economic impacts. According to the MegaScale (Jiang et al., 2024) report, the average recovery time from an interruption is about 15 minutes. Considering the Llama-3  [^1] https://ai.meta.com/blog/meta-llama-3/ model, which utilizes 16,000 NVIDIA H100 GPUs priced between \$2.29 to \$2.49 per hour  [^2] https://lambdalabs.com/, each interruption due to a Loss Spike can lead to direct economic losses ranging from approximately \$9,160 to \$9,960. Our proposed adaptive gradient clipping method effectively reduces the frequency of Loss Spikes, thereby enhancing training stability and minimizing economic waste.
6.2 Limitations
Resource constraints prevented this study from conducting long-term training experiments on the Llama-2 open-source model. However, our internal business validations indicate that the method remains effective throughout the mid to late stages of training. Future plans, resources permitting, include conducting comprehensive experiments on the Llama-2 model and sharing the findings with the academic community.
7 Conclusion
This paper systematically addresses the critical challenge of loss spikes in large model pre-training through gradient dynamics analysis. Our key insight reveals two limitations in conventional approaches: temporal gradient norm decay requires adaptive thresholds, while parameter-specific gradient heterogeneity demands localized control. These findings motivate AdaGC, which integrates exponential moving average-based threshold adaptation with per-parameter clipping while maintaining Adam-comparable $\mathcal{O}(1/\sqrt{T})$ convergence rate - proving local clipping preserves theoretical convergence properties. Comprehensive experiments demonstrate AdaGC’s dual benefits: complete loss spike elimination across LLMs (3.5% WikiText perplexity reduction on Llama-2 7B) and LVLMs (25% faster CLIP convergence than StableAdamW), coupled with enhanced training stability and final model quality. The consistent success across architectures and optimizers validates that jointly addressing temporal adaptation and spatial parameter characteristics provides an effective stabilization paradigm for large-scale training.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Convergence Proof
In this section, we provide the necessary assumptions and lemmas for the proofs of Theorem 4.1.
Notations
The $k$-th component of a vector $v_{t}$ is denoted as ${v}_{t,k}$. Other than that, all computations that involve vectors shall be understood in the component-wise way. We say a vector $v_{t}\geq 0$ if every component of $v_{t}$ is non-negative, and $v_{t}\geq w_{t}$ if $v_{t,k}\geq w_{t,k}$ for all $k=1,2,\ldots,d$. The $\ell_{1}$ norm of a vector $v_{t}$ is defined as $\|v_{t}\|_{1}=\sum_{k=1}^{d}|{v}_{t,k}|$. The $\ell_{2}$ norm is defined as $\|v_{t}\|^{2}=\langle v_{t},v_{t}\rangle=\sum_{k=1}^{d}|{v}_{t,k}|^{2}$. Given a positive vector $\hat{\eta}_{t}$, it will be helpful to define the following weighted norm: $\|v_{t}\|^{2}_{\eta_{t}}=\langle v_{t},\hat{\eta}_{t}v_{t}\rangle=\sum_{k=1}^{d}\hat{\eta}_{t,k}|{v}_{t,k}|^{2}$.
Assumption A.1.
The function f is lower bounded by $\underline{f}$ with L-Lipschitz gradient.
Assumption A.2.
The gradient estimator $g$ is unbiased with bounded norm, e.g,
$\begin{split}&\mathbb{E}[g|x_{t}]=\nabla f(x_{t}),\ \|g_{t}\|\leq G.\end{split}$
Assumption A.3.
The coefficient of clipping $h_{t,i}$ is lower bounded by some $h_{0}>0$.
Assumption A.4.
$\|g_{t}-\nabla f(x_{t})\|\leq p\|\nabla f(x_{t})\|$ holds for some $p<1$ and for all t.
Remark A.5.
Assumption A.1 and Assumption A.2 are widely used in the proof of optimization algorithm with adaptive learning rates (Reddi et al., 2018). Assumption A.3 is because the gradient norm changes slowly when training the neural network, and the last assumption holds when the batch size is large enough.
Lemma A.6.
Let $\zeta:=\beta_{1}^{2}/{\beta_{2}}$. We have the following estimate
$m_{t}^{2}\leq\frac{1}{(1-\zeta)(1-\beta_{2})}v_{t},~{}\forall t.$ (4)
Proof.
By the iteration formula $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\hat{g}_{t}$ and $m_{0}=0$, we have
$m=\sum_{i=1}^{t}\beta_{1}^{t-i}(1-\beta_{1})\hat{g}_{i}.$
Similarly, by $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\hat{g}_{t}^{2}$ and $v_{0}=0$, we have
$v_{t}=\sum_{i=1}^{t}\beta_{2}^{t-i}(1-\beta_{2})\hat{g}_{i}^{2}$
It follows by arithmetic inequality that
$\begin{split}m_{t}^{2}&=\left(\sum_{i=1}^{t}\frac{(1-\beta_{1})\beta_{1}^{t-i}}{\sqrt{(1-\beta_{2})\beta_{2}^{t-i}}}\sqrt{(1-\beta_{2})\beta_{2}^{t-i}}\hat{g}_{i}\right)^{2}\\ &\leq\left(\sum_{i=1}^{t}\frac{(1-\beta_{1})^{2}\beta_{1}^{2(t-i)}}{(1-\beta_{2})\beta_{2}^{t-i}}\right)\left(\sum_{i=1}^{t}(1-\beta_{2})\beta_{2}^{t-i}\hat{g}_{i}^{2}\right)=\left(\sum_{i=1}^{t}\frac{(1-\beta_{1})^{2}\beta_{1}^{2(t-i)}}{(1-\beta_{2})\beta_{2}^{t-i}}\right)v_{t}.\end{split}$
Further, we have
$\sum_{i=1}^{t}\frac{(1-\beta_{1})^{2}\beta_{1}^{2(t-i)}}{(1-\beta_{2})\beta_{2}^{t-i}}\leq\frac{1}{1-\beta_{2}}\sum_{i=1}^{t}\left(\frac{\beta_{1}^{2}}{\beta_{2}}\right)^{t-i}=\frac{1}{1-\beta_{2}}\sum_{k=0}^{t-1}\zeta^{k}\leq\frac{1}{(1-\zeta)(1-\beta_{2})}.$
The proof is completed. ∎
Lemma A.7.
The following estimate holds
$\sum_{t=1}^{T}\|\Delta_{t}\|^{2}\leq\frac{\alpha^{2}G^{2}}{\epsilon}$
Proof.
By using the definition of $m_{t}$, it holds $\|m_{t}\|^{2}\leq G^{2}$.
Then, $\|\Delta_{t}\|^{2}=\|\frac{\alpha_{t}m_{t}}{\sqrt{v_{t}}+\epsilon}\|^{2}\leq\frac{G^{2}}{\epsilon}\alpha_{t}^{2}$ by using the definition of $\Delta_{t}$.
Therefore, $\small\sum_{t=1}^{T}\|\Delta_{t}\|^{2}\leq\frac{G^{2}}{\epsilon}\sum_{t=1}^{T}\frac{\alpha^{2}}{T}=\frac{G^{2}\alpha^{2}}{\epsilon}$.
Lemma A.8.
With the Assumption A.3 and A.4, it holds that
$\mathbb{E}\left\langle\nabla f\left(\theta_{t}\right),\hat{\eta}_{t}\hat{g}_{t}\right\rangle\geq h_{0}\mathbb{E}\left\|\nabla f\left(\theta_{t}\right)\right\|_{\hat{\eta}_{t}}^{2}.$
Proof.
According to Assumption A.4, it holds that
$\begin{split}\left\langle\nabla_{i}f\left(\theta_{t}\right),g_{t,i}\right\rangle&=-\frac{1}{2}\left(\left\|\nabla_{i}f\left(\theta_{t}\right)-g_{t,i}\right\|^{2}-\left\|\nabla_{i}f\left(\theta_{t}\right)\right\|^{2}-\left\|g_{t,i}\right\|^{2}\right)\\ &\geq(1-p^{2})\left\|\nabla_{i}f\left(\theta_{t}\right)\right\|^{2}\geq 0.\end{split}$
Thus, it holds that
$\begin{split}\mathbb{E}\left[\left\langle\nabla f(x_{t}),\hat{\eta}_{t}\hat{g}_{t}\right\rangle\right]&=\mathbb{E}\left[\sum_{i}\left\langle\nabla_{i}f(\theta_{t}),h_{t,i}\hat{\eta}_{t,i}g_{t,i}\right\rangle\right]\\ &\geq h_{0}\mathbb{E}\left[\sum_{i}\left\langle\nabla_{i}f(x_{t}),h_{t,i}\hat{\eta}_{t,i}g_{t,i}\right\rangle\right]\\ &=h_{0}\mathbb{E}\left\langle\nabla f(\theta_{t}),\hat{\eta}_{t}g_{t}\right\rangle=h_{0}\mathbb{E}\|\nabla f(\theta_{t})\|^{2}_{\hat{\eta}_{t}}.\end{split}$
Let $\Delta_{t}:=\theta_{t+1}-\theta_{t}=-\alpha_{t}m_{t}/(\sqrt{v_{t}}+\epsilon)$. Let $\hat{v}_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\delta_{t}^{2}$, where $\delta_{t}^{2}=\mathbb{E}_{t}\left[\hat{g}_{t}^{2}\right]$ and let $\hat{\eta}_{t}=\alpha_{t}/\sqrt{\hat{v}_{t}+\epsilon}$.
Lemma A.9.
Let $M_{t}=\mathbb{E}\left[\langle\nabla f(\theta_{t}),\Delta_{t}\rangle+L\|\Delta_{t}\|^{2}\right]$. Let $\alpha_{t}=\alpha/\sqrt{T}$ and $\beta_{2}=1-\beta/T$. Then, for $T\geq 1$ we have
$\begin{split}\sum_{t=1}^{T}M_{t}\leq\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2},\end{split}$ (5)
where $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right).$
Proof.
To split $M_{t}$, firstly we introduce the following two equalities. Using the definitions of $v_{t}$ and $\hat{v}_{t}$, we obtain
$\begin{split}\frac{\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}}{\sqrt{v_{t}}+\epsilon}&=\frac{\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}+\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}\left(\frac{1}{\sqrt{v_{t}}+\epsilon}-\frac{1}{\sqrt{\hat{v}_{t}}+\epsilon}\right)\\ &=\left(1-\beta_{1}\right)\hat{\eta}_{t}\hat{g}_{t}+\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}\frac{\left(1-\beta_{2}\right)\left(\sigma_{t}^{2}-\hat{g}_{t}^{2}\right)}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{\hat{v}_{t}}+\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}}\right)}\\ &=\left(1-\beta_{1}\right)\hat{\eta}_{t}\hat{g}_{t}+\left(1-\beta_{1}\right)\hat{\eta}_{t}\hat{g}_{t}\frac{\left(1-\beta_{2}\right)\left(\sigma_{t}^{2}-\hat{g}_{t}^{2}\right)}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}}\right)}\end{split}$
In addition, we can obtain:
$\begin{split}&\beta_{1}\alpha_{t}m_{t-1}\left(\frac{1}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &=\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\sqrt{\beta_{2}}\right)\epsilon}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)}\\ &=\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\hat{v}_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\\ &\qquad+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\left(\frac{1}{\sqrt{\hat{v}_{t}}+\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &\qquad+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\sqrt{\beta_{2}}\right)\epsilon}{\left(\sqrt{\hat{v}_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)}+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\sqrt{\beta_{2}}\right)\epsilon}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}\left(\frac{1}{\sqrt{\hat{v}_{t}}+\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &=\beta_{1}m_{t-1}\hat{\eta}_{t}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\\ &\qquad+\beta_{1}\hat{\eta}_{t}m_{t-1}\frac{(1-\beta_{2})^{2}\hat{g}_{t}^{2}(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}})}\\ &\qquad+\beta_{1}\hat{\eta}_{t}m_{t-1}\frac{(1-\sqrt{\beta_{2}})\epsilon}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}+\beta_{1}\hat{\eta}_{t}m_{t-1}\frac{(1-\sqrt{\beta_{2}})(1-\beta_{2})\epsilon(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)}.\end{split}$
For simplicity, we denote
$\begin{split}&A_{t}^{1}=\left(1-\beta_{1}\right)\sqrt{\hat{\eta}_{t}}\hat{g}_{t}\frac{\left(1-\beta_{2}\right)\left(\sigma_{t}^{2}-\hat{g}_{t}^{2}\right)}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}}\right)}\\ &A_{t}^{2}=\beta_{1}m_{t-1}\sqrt{\hat{\eta}_{t}}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\\ &A_{t}^{3}=\beta_{1}\sqrt{\hat{\eta}_{t}}m_{t-1}\frac{(1-\beta_{2})^{2}\hat{g}_{t}^{2}(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}})}\\ &A_{t}^{4}=\beta_{1}\sqrt{\hat{\eta}_{t}}m_{t-1}\frac{(1-\sqrt{\beta_{2}})\epsilon}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}\\ &A_{t}^{5}=\beta_{1}\sqrt{\hat{\eta}_{t}}m_{t-1}\frac{(1-\sqrt{\beta_{2}})(1-\beta_{2})\epsilon(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)}\\ \end{split}$
Then, we obtain
$\begin{split}\Delta_{t}-\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\Delta_{t-1}&=-\frac{\alpha_{t}m_{t}}{\sqrt{v_{t}}+\epsilon}+\frac{\beta_{1}\alpha_{t}m_{t-1}}{\sqrt{\beta_{2}v_{t-1}+\sqrt{\beta_{2}}\epsilon}}\\ &=-\frac{(1-\beta_{1})\alpha_{t}\hat{g}_{t}}{\sqrt{v_{t}}+\epsilon}+\beta_{1}\alpha_{t}m_{t-1}\left(\frac{1}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &=-(1-\beta_{1})\hat{\eta}_{t}\hat{g}_{t}-\sqrt{\hat{\eta}_{t}}A_{t}^{1}+\sqrt{\hat{\eta}_{t}}A_{t}^{2}+\sqrt{\hat{\eta}_{t}}A_{t}^{3}+\sqrt{\hat{\eta}_{t}}A_{t}^{4}+\sqrt{\hat{\eta}_{t}}A_{t}^{5}\end{split}$
Thus, it holds that
$\begin{split}\mathbb{E}\left\langle\nabla f(\theta_{t}),\Delta_{t}\right\rangle&=\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left\langle\nabla f(\theta_{t}),\Delta_{t-1}\right\rangle+\mathbb{E}\left\langle\nabla f(\theta_{t}),\Delta_{t}-\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\Delta_{t-1}\right\rangle\\ &=\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+\mathbb{E}\langle\nabla f(\theta_{t})-\nabla f(\theta_{t-1}),\Delta_{t-1}\rangle\right)\\ &\qquad+\mathbb{E}\langle\nabla f(\theta_{t}),-(1-\beta_{1})\hat{\eta}_{t}\hat{g}_{t}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),-\sqrt{\hat{\eta}_{t}}A_{t}^{1}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{2}\rangle\\ &\qquad+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{3}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{4}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{5}\rangle\end{split}$ (6)
For the first term of (6), it holds that
$\begin{split}&\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+\mathbb{E}\langle\nabla f(\theta_{t})-\nabla f(\theta_{t-1}),\Delta_{t-1}\rangle\right)\\ &\leq\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+\mathbb{E}\|\nabla f(\theta_{t})-\nabla f(\theta_{t-1})\|\|\Delta_{t-1}\|\right)\\ &\leq\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+L\mathbb{E}\|\Delta_{t-1}\|^{2}\right)\\ &=\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}M_{t-1}\end{split}$
For the second term of (6), it holds that
$\mathbb{E}\langle\nabla f(\theta_{t}),-(1-\beta_{1})\hat{\eta}_{t}\hat{g}_{t}\rangle\leq-(1-\beta_{1})h_{0}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}.$
For the rest of the terms, it holds that
$\begin{split}\mathbb{E}\langle\nabla f(\theta_{t}),-\sqrt{\hat{\eta}_{t}}A_{t}^{1}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{1}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{2}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{2}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{3}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{3}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{4}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{4}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{5}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{5}\right\|^{2}\\ \end{split}$
On the other hand, it holds that
$\begin{split}&\left\|A_{t}^{1}\right\|^{2}\leq(1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{T\epsilon^{3}},\left\|A_{t}^{2}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{4}}{T\beta_{2}\epsilon^{3}},\left\|A_{t}^{3}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{2}}{(1-\zeta)\epsilon T\beta_{2}},\\ &\left\|A_{t}^{4}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{4}}{T\beta_{2}},\left\|A_{t}^{5}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{2}}{(1-\zeta)\beta_{2}T}\end{split}$
Define $N_{t}=\frac{C_{2}}{T}+L\mathbb{E}\|\Delta_{t}\|^{2}$, where $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right)$. It holds that
$M_{t}\leq\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}M_{t-1}+N_{t}-\frac{1-\beta_{1}}{2}\hat{\eta}_{t}\mathbb{E}\|\nabla f(\theta_{t})\|^{2}_{\hat{\eta}_{t}}\leq\sum_{i=1}^{t}\sqrt{\zeta}^{t-i}N_{i}-\frac{1-\beta_{1}}{2}h_{0}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}$
Thus, by summing $t$ from 1 to $T$, it holds that
$\begin{split}\sum_{t=1}^{T}M_{t}&\leq\sum_{t=1}^{T}\sum_{i=1}^{t}\sqrt{\zeta}^{t-i}N_{i}-\frac{(1-\beta_{1})h_{0}}{2}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\\ &\leq\frac{1}{1-\sqrt{\zeta}}\sum_{t=1}^{T}N_{t}-\frac{(1-\beta_{1})h_{0}}{2}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\\ &\leq\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}.\end{split}$
Lemma A.10.
Let $\tau$ be randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities $p_{\tau}=\frac{1}{T}$. We have the following estimate:
$\mathbb{E}[\|\nabla f\left(\theta_{\tau}\right)\|^{2}]\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f\left(\theta_{t}\right)\|_{\hat{\eta_{t}}}^{2}\right].$
Proof.
Note that $\|\hat{v}_{t}\|_{1}=\beta_{2}\|v_{t-1}\|_{1}+\left(1-\beta_{2}\right)\|\sigma_{t}\|^{2}$ and $\|\hat{g}_{t}\|\leq G$. It is straightforward to prove $\|v_{t}\|_{1}\leq G^{2}$. Hence, we have $\|\hat{v}_{t}+\epsilon\|_{1}\leq G^{2}+\epsilon d$.
Utilizing this inequality, we have
$\begin{split}\|\nabla f\left(\theta_{t}\right)\|^{2}&=\frac{\|\nabla f\left(\theta_{t}\right)\|^{2}}{\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}}\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}=\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}\sum_{k=1}^{d}\frac{|\nabla_{k}f\left(\theta_{t}\right)|^{2}}{\sqrt{\sum_{l=1}^{d}\hat{v}_{t,l}+\epsilon}}\\ &\leq\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}\alpha_{t}^{-1}\sum_{k=1}^{d}\frac{\alpha_{t}}{\sqrt{\hat{v}_{t,k}+\epsilon}}|\nabla_{k}f\left(\theta_{t}\right)|^{2}=\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}\alpha_{t}^{-1}\|\nabla f\left(\theta_{t}\right)\|_{\hat{\eta}_{t}}^{2}\\ &\leq\sqrt{G^{2}+\epsilon d}\alpha_{t}^{-1}\|\nabla f\left(\theta_{t}\right)\|_{\hat{\eta}_{t}}^{2}\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha_{T}}\|\nabla f\left(\theta_{t}\right)\|^{2}_{\hat{\eta}_{t}}.\end{split}$
Then, by using the definition of $\theta_{\tau}$, we obtain
$\mathbb{E}\left[\|\nabla f\left(\theta_{\tau}\right)\|^{2}\right]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\|\nabla f\left(\theta_{t}\right)\|^{2}\right]\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f\left(\theta_{t}\right)\|^{2}_{\hat{\eta}_{t}}\right].$
Thus, the desired result is obtained. ∎
Theorem A.11.
Let $\{\theta_{t}\}$ be a sequence generated by AdaGC for initial values $\theta_{1}$ and $m_{0}=v_{0}=0$. Assumptions A.1 to A.4 hold. With the hyperparameters $\alpha_{t}=\alpha/\sqrt{T}$,$\beta_{2}=1-\beta/T$ and $\zeta=\beta_{1}^{2}/\beta_{2}<1$. Let $\tau$ be randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities. We have
$\mathbb{E}\|\nabla f(\theta_{\tau})\|^{2}\leq\frac{C}{\sqrt{T}}$
where $C=\frac{\sqrt{G^{2}+\epsilon d}}{\alpha}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)$ and $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right)$
Proof.
With the Lipschitz continuity condition of $f$, it holds that
$\mathbb{E}f(\theta_{t+1})\leq\mathbb{E}\left[f(\theta_{t})+\langle\nabla f(\theta_{t}),\Delta_{t}\rangle+\frac{L}{2}\|\Delta_{t}\|^{2}\right]\leq\mathbb{E}f(\theta_{t})+M_{t}.$
By summing $t$ from $1$ to $T$, it holds that
$\mathbb{E}f(\theta_{T+1})\leq f(\theta_{1})+\sum_{t=1}^{T}M_{t}\leq f(\theta_{1})+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}$
Thus, it holds that
$\begin{split}\mathbb{E}\left[\|\nabla f(\theta_{\tau}\|^{2}\right]&\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\right]\\ &\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\left(f(\theta_{1})-\mathbb{E}[f(\theta_{T+1})]+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)\\ \ &\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)\end{split}$
Appendix B LLMs Hpyer-Parameters
Table 2 shows the hyper-parameters used in our experiments on LLMs. Note that Llama-2 Tiny is a smaller version of the Llama-2 model, designed based on the network configuration of the GPT-2 model with 345 million parameters. We named it Llama-2 Tiny to facilitate our initial experimental exploration.
Table 2: Hyper-parameters used in our LLMs experiments. $\lambda_{abs}$ represents the absolute global clipping threshold of GlobalGC. $\lambda_{rel}$ and $\beta$ represent the relative clipping threshold and the momentum of our AdaGC, respectively.
Model LLaMA-tiny LLaMA-7B LLaMA-13B GPT-2
Precision bfloat16 bfloat16 bfloat16 bfloat16
Layer num 24 32 40 24
Hidden dim size 1024 4096 5120 1024
FFN dim size 2371 11008 13824 4096
Attention heads 16 32 40 16
Sequence length 2048 2048 2048 2048
Batch size 512 2048 2048 512
iterations 36000 9000 9000 30000
Learning rate $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$
Lr decay style cosine cosine cosine cosine
Warmup iterations 2000 2000 2000 300
Weight decay 0.1 0.1 0.1 0.1
$\lambda_{abs}$ 1.0 1.0 1.0 1.0
$\lambda_{rel}$ 1.05 1.05 1.05 1.05
$\beta$ 0.98 0.98 0.98 0.98
Appendix C Experimental Details for CLIP
In addition to pre-training tasks for Large Language Models (LLMs), we have conducted pre-training for large-scale Vision-Language Models (LVLMs), with a specific focus on the widely acknowledged CLIP model (Radford et al., 2021). We train the CLIP models on the LAION-400M (Schuhmann et al., 2021) dataset and conduct zero-shot evaluations on the ImageNet (Russakovsky et al., 2015) dataset. We opt for the ViT-Base configuration, which contains 151 million parameters. The ViT-Base model is trained over 20K steps, observing 320M samples. In alignment with (Wortsman et al., 2023a), we implement patch-dropout 0.5 (Li et al., 2023), learning rate 0.002 and weight decay 0.2, with the initial 5K steps dedicated to linear warmup (Goyal et al., 2017) and the subsequent steps follow a cosine decay pattern (Loshchilov & Hutter, 2016).
Appendix D Results of Batch Size Scalability Experiments
(a) batch size = $512$, lr = $3.0\times 10^{-4}$
(b) batch size = $2048$, lr = $6.0\times 10^{-4}$
(c) batch size = $4096$, lr = $8.5\times 10^{-4}$
(d) batch size = $8192$, lr = $1.2\times 10^{-3}$
Figure 9: AdaGC’s large-batch scalability on GPT-2 345M: Maintains 0 loss spikes (vs GlobalGC’s 100% rate) with batch sizes up to 8192 (16$\times$ baseline) and proportional learning rate scaling, achieving 1.28%-19.55% perplexity reduction.
(a) batch size = $512$, lr = $3.0\times 10^{-4}$
(b) batch size = $2048$, lr = $6.0\times 10^{-4}$
(c) batch size = $4096$, lr = $8.5\times 10^{-4}$
(d) batch size = $8192$, lr = $1.2\times 10^{-3}$
Figure 10: AdaGC’s large-batch scalability on Llama-2 Tiny: Cross-architecture validation shows 6.19% lower perplexity at 8192 batch size.

### v2

@@
-AdaGC: Improving Training Stability for Large Language Model Pretraining
+AdaGC: Enhancing LLM Pretraining Stability via Adaptive Gradient Clipping
 Abstract
-Large Language Models (LLMs) face increasing loss spikes during scaling, undermining training stability and final performance.
-While gradient clipping mitigates this issue, traditional global approaches poorly handle parameter-specific gradient variations and decaying gradient norms.
-We propose AdaGC, an adaptive gradient clipping framework that automatically adjusts local thresholds per parameter through exponential moving average of gradient norms.
-Theoretical analysis proves AdaGC’s convergence under non-convex conditions.
-Extensive experiments demonstrate significant improvements: On Llama-2 7B/13B, AdaGC completely eliminates loss spikes while reducing WikiText perplexity by 3.5% (+0.14pp LAMBADA accuracy) for 7B and achieving 0.65% lower training loss with 1.47% reduced validation perplexity for 13B compared to global clipping.
-For CLIP ViT-Base, AdaGC converges 25% faster than StableAdamW with full spike elimination.
-The method shows universal effectiveness across architectures (Llama-2 7B/13B) and modalities (CLIP), with successful integration into diverse optimizers like AdamW and Lion.
-Source code will be released on GitHub.
+Loss spikes remain a persistent obstacle in large-scale language model pretraining.
+While previous research has attempted to identify the root cause of loss spikes by investigating individual factors, we observe that, in practice, such spikes are typically triggered by the confluence of heterogeneous factors.
+Empirically, loss spikes may arise from a combination of data outliers, hardware or transient computational faults, numerical precision issues, and hyperparameter settings.
+Regardless of the underlying cause, these spikes manifest as unstable optimizer updates, as abnormal gradients contaminate both first- and second-moment states.
+In this paper, we propose a principled gradient-centric remedy: AdaGC, an adaptive per-tensor gradient clipping scheme that mitigates such contamination by bounding gradient norms relative to a tensor-wise exponential moving average of their historical clipped values.
+AdaGC is optimizer-agnostic, introduces negligible memory overhead, and reduces communication costs compared to GlobalGC, particularly in hybrid-parallel distributed training environments.
+Experiments on Llama-2 7B, Mixtral 8×1B, and ERNIE 10B-A1.4B demonstrate that AdaGC robustly eliminates training instabilities, consistently reducing spike scores to zero for all models and improving downstream accuracy over GlobalGC by 1.32%, 1.27%, and 2.48%, respectively.
+Furthermore, AdaGC seamlessly integrates with optimizers such as Muon and Lion, consistently yielding higher average accuracy and zero spike scores.
 Machine Learning, ICML
 1 Introduction
-Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b) have become a cornerstone of modern Natural Language Processing (NLP) research, demonstrating outstanding performance in a wide range of language tasks.
-These models, characterized by their extensive parameter counts and vast training datasets, are designed to capture the subtleties of human language.
-However, as these models grow in size and complexity, they often encounter a significant training obstacle: the loss spike—a sharp, unexpected increase in training loss that can seriously undermine the training process’s efficiency and effectiveness (Chowdhery et al., 2023; Scao et al., 2022; Takase et al., 2023; Zeng et al., 2022).
-Among the various methods for addressing spike issues, gradient clipping is an effective solution.
-Existing gradient clipping approaches (Pascanu et al., 2013) typically apply a predefined threshold to the gradient norm of the entire neural network.
-However, we have identified two significant shortcomings with this global threshold method.
-First, as the gradient magnitudes decrease during training, an excessively small threshold throughout the process can slow down the training speed, while a consistently large threshold fails to effectively mitigate spike issues in the later stages of training.
-Second, due to the inherent differences between structures in neural networks (Zhang et al., 2024), particularly in transformers, applying a uniform clipping coefficient across all layers does not adequately resolve the problems caused by spikes, as illustrated in the Figure 1.
-To address these two issues with global gradient clipping, we propose using local gradient clipping along with an adaptive update of the clipping threshold for each parameter.
-Specifically, we introduce a novel method called Adaptive Gradient Clipping based on Local Gradient Norm (AdaGC).
-AdaGC is specifically designed to mitigate loss spike by selectively suppressing abnormal gradients, thereby maintaining training stability and enhancing model performance.
-At its core, AdaGC employs the Exponential Moving Average (EMA) to accurately track historical gradient norm variations at the tensor level.
-This enables the precise identification and adaptive clipping of anomalous local gradients, effectively containing the loss spike.
-Convergence analysis shows AdaGC maintains the $\mathcal{O}(1/\sqrt{T})$ rate comparable to Adam in non-convex settings, confirming that localized clipping preserves convergence properties.
-Comprehensive experiments on Llama-2 7B/13B (Touvron et al., 2023b) and CLIP ViT-Base (Radford et al., 2021) demonstrate complete loss spike elimination while improving model convergence (3.5% perplexity reduction, 25% faster convergence than StableAdamW (Wortsman et al., 2023a)), with successful integration into modern optimizers like AdamW (Loshchilov & Hutter, 2017) and Lion (Chen et al., 2024).
-In summary, our main contributions are three-fold:
-  • Identification of two critical limitations in global gradient clipping methods: inadequate handling of decaying gradient norms and parameter-wise heterogeneity, motivating the need for localized adaptive clipping.
-  • Proposal of Adaptive Gradient Clipping (AdaGC), featuring parameter-wise threshold adaptation through exponential moving averages, effectively addressing both temporal gradient decay and spatial parameter variations.
-  • Comprehensive empirical validation across architectures (Llama-2 7B/13B) and modalities (CLIP ViT-Base), demonstrating complete loss spike elimination with measurable performance gains, plus convergence analysis confirming comparable rates to standard optimizers.
+(a) Mixtral
+(b) ERNIE
+(c) ERNIE
+(d) ERNIE
+Figure 1: Reproduced cases of loss spikes and mitigation via resuming.
+Loss spikes are triggered by (a) increasing $\beta_{2}$ or (b) reducing $\epsilon$ in AdamW, (c) using lower-precision RMSNorm, even under global gradient clipping, and (d) are resolved by resuming due to stochasticity in FlashAttention backward passes.
+Table 1: Common sources of loss spikes in LLM pretraining, with typical triggers and practical mitigations.
+Source Typical trigger Common mitigation
+Data quality Noisy/ill-formed samples Resume and skip data batches
+Hardware / transient faults GPU silent errors Fault tolerance and machine replacement
+Numerical precision Activation outlier BF16 to FP32 or blockwise quantization
+Optimizer / LN hyperparams AdamW/RMSNorm $\epsilon$; AdamW $\beta_{2}$ Careful hyperparameter selection
+The rapid scaling of large language models (LLM) has introduced new challenges in pretraining stability, often manifesting as abrupt loss spikes or transient divergences across a wide range of model architectures and data scales (Chowdhery et al., 2023; Touvron et al., 2023; Liu et al., 2024; Team et al., 2025; Baidu-ERNIE-Team, 2025).
+Despite extensive empirical studies, the fundamental causes of these instabilities remain elusive.
+Recent research, alongside our own analyses, indicates that loss spikes can arise from a variety of sources, including: (i) data quality issues (Chowdhery et al., 2023); (ii) hardware or transient computational faults (Su, 2025); (iii) variations in numerical precision (for example, FP32 typically offers greater robustness than BF16, whereas FP8 can sometimes enhance stability by suppressing outlier values via implicit quantization (Han, 2024; Liu et al., 2024)); and (iv) the selection of optimizer and layer normalization hyperparameters, such as the $\epsilon$ parameter in RMSNorm or AdamW, and $\beta_{2}$ in AdamW (Ma et al., 2021; Cattaneo & Shigida, 2025; Bai et al., 2025).
+For instance, we observe that increasing $\beta_{2}$ or decreasing $\epsilon$ in AdamW can trigger loss spikes, whereas increasing the precision of RMSNorm from BF16 to FP32 significantly improves stability.
+In summary, Table 1 presents the common sources, typical triggers, and practical mitigations for loss spikes identified from both previous studies and our experimental findings, while Figure 1 demonstrates several representative cases we have reproduced.
+Although the upstream causes of instability are diverse and often subtle, these events consistently converge at the optimizer level, manifesting as abnormal gradients.
+Such outlier gradients are incorporated into the optimizer’s first- and second-moment estimates, thereby corrupting parameter updates and propagating instability through subsequent training.
+Notably, we find that even resuming interrupted training (while keeping the random seed and data unchanged) can mitigate a loss spike, merely due to the stochastic nature of $dQ$, dKdK, and $dV$ in FlashAttention (Dao, 2023) (see Figure 1(d)).
+This observation further suggests that, in certain model states, even minute numerical differences can trigger a loss spike, with gradient outliers playing a critical role in both the initiation and propagation of instabilities during optimizer state updates.
+While the slight stochasticity introduced by FlashAttention can sometimes circumvent a loss spike, repeatedly interrupting and resuming training imposes substantial computational overhead.
+Given that these instabilities stem from diverse upstream causes but ultimately converge at the optimizer level, our work does not attempt to identify the precise root causes.
+Instead, we adopt a gradient-centric perspective: irrespective of the initial trigger, loss spikes consistently arise when outlier gradients contaminate the optimizer states.
+Therefore, by preventing such gradients from entering the first- and second-moment accumulators, we provide a unified and effective strategy to mitigate training instability.
+A standard mitigation strategy is global gradient clipping (GlobalGC), which bounds the global $\ell_{2}$ norm of the aggregated gradient.
+However, this approach is fundamentally mismatched to modern large-scale pretraining in two key respects: (1) Temporal mismatch: The optimal global clipping threshold typically decreases over the course of training; a fixed threshold risks under-clipping in later phases.
+(2) Spatial mismatch: Gradient statistics and rare spikes vary asynchronously across different parameter tensors, making a single global threshold insufficient—protecting one tensor may under-serve or over-constrain others.
+To address these challenges, we introduce Adaptive Gradient Clipping based on Local Gradient Norm (AdaGC): a simple, per-tensor clipping rule that leverages an EMA of each tensor’s historical gradient norm as a reference.
+Each tensor’s gradient is clipped relative to its own EMA, preventing transient outliers from contaminating the first- and second-moment accumulators and, ultimately, the parameter updates.
+A brief warm-up period applies global clipping and initializes the EMA to avoid early overestimation.
+AdaGC is optimizer-agnostic and can be seamlessly integrated with AdamW, Lion, and Muon.
+Our main contributions are as follows:
+  • A unified, gradient-centric perspective: We clarify how loss spikes universally propagate via abnormal gradients polluting optimizer states, irrespective of their origin, motivating intervention at the gradient level prior to moving-average accumulation.
+  • An adaptive, per-tensor clipping rule: By tracking each tensor’s gradient norm statistics with an EMA, AdaGC provides both temporal adaptivity and spatial specificity, suppressing outliers while minimally disturbing typical learning dynamics.
+  • System efficiency: We analyze computational and communication overhead, showing that AdaGC reduces communication relative to GlobalGC under hybrid parallel distributed training.
+[…]
 2 Related Work
-Loss spikes present a significant challenge when scaling models, potentially impeding learning or destabilizing the training process.
-Various strategies have been proposed to enhance the stability of pre-training in large language models.
-Training Strategies: Effective training strategies are pivotal for achieving stability.
-Adopting the bfloat16 format (Wang & Kanwar, 2019) is a cornerstone for stable training.
-Using smaller learning rates (Wortsman et al., 2023b; Zhang et al., 2022) also promotes stable training dynamics.
-To mitigate loss spikes, PaLM (Chowdhery et al., 2023) resumes training from a checkpoint approximately 100 steps before the spike and bypasses 200-500 data batches to avoid those preceding and during the spike.
-Varying the sequence length (Li et al., 2022) is also effective for stabilizing the pre-training of large language models.
-Model Architecture: Advanced architectures are crucial for stability in training large-scale models.
-Research shows that Pre-LN Transformers offer enhanced stability over Post-LN Transformers (Xiong et al., 2020; Vaswani et al., 2017; Liu et al., 2020), both theoretically and practically.
-Thus, contemporary studies predominantly employ Pre-LN Transformers for building large-scale models.
-Techniques such as NormHead (Yang et al., 2023), which normalizes output embeddings, and RMSNorm (Zhang & Sennrich, 2019), as reported in Llama (Touvron et al., 2023a), contribute to stable training.
-Applying layer normalization to the embedding layer (Scao et al., 2022), normalizing the input vector before the standard softmax function (NormSoftmax) (Jiang et al., 2023), and the shrink embedding gradient technique (Zeng et al., 2022) are also beneficial.
-Model Initialization: Proper model initialization significantly enhances training stability.
-Scaling the embedding layer with an appropriate constant (Takase et al., 2023) reinforces LLM pre-training stability.
-Initializing Transformers with small parameter values (Nguyen & Salazar, 2019) also improves pre-training stability.
-Some researchers propose that removing layer normalizations in Transformers is feasible with specific initialization methods (Zhang et al., 2019; Huang et al., 2020).
-Auxiliary Loss Function: Incorporating an auxiliary loss function alongside conventional language modeling loss enhances training stability.
-Notable examples include Max-z loss, which benefits the pre-training process (Chowdhery et al., 2023; Wortsman et al., 2023b; Yang et al., 2023).
-Gradient/Update Clipping: Gradient and update clipping achieve stability by limiting the magnitude of parameter updates, preventing excessively large weight updates.
-Global gradient clipping (Pascanu et al., 2013) is prevalent, with innovative approaches like adaptive gradient clipping (Brock et al., 2021) and Clippy (Tang et al., 2023), which use model weights to adjust the clipping threshold.
-Alternatives like Adafactor (Shazeer & Stern, 2018), StableAdamW (Wortsman et al., 2023a), and LAMB (You et al., 2019) offer update clipping techniques better suited for stability training of large-scale models.
+Table 2: Comparison of major gradient/update clipping methods for stability.
+Here, $\boldsymbol{\theta}_{t}$ denotes the model parameters, $\boldsymbol{g}_{t}$ the gradients, $\Delta_{t}$ the optimizer update, $\eta_{t}$ the learning rate, $\lambda_{abs}$ the absolute threshold, and $\lambda_{rel}$ the relative threshold.
+Method Algorithm Gradient Update Granularity Threshold Type
+GlobalGC (Pascanu et al., 2013) $\min\{1.0,\lambda_{abs}\frac{1}{\|\boldsymbol{g}_{t}\|}\}$ Global Fixed constant
+ClipByValue $clamp(-\lambda_{abs},\lambda_{abs})$ Element Fixed constant
+AGC (Brock et al., 2021) $\min\{1.0,\lambda_{rel}\frac{\|\boldsymbol{\theta}_{t}\|}{\|\Delta_{t}\|}\}$ Unit Weight $\ell_{2}$ norm
+Clippy (Tang et al., 2023) $\min\{1.0,\min(\frac{\lambda_{rel}\|\boldsymbol{\theta}_{t}\|_{\infty}+\lambda_{abs}}{\eta_{t}*\|\Delta_{t}\|_{\infty}})\}$ Tensor Weight $\ell_{\infty}$ norm
+AdaGC (ours) $\min\{1.0,\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|}\}$ Tensor EMA of gradient norm
+$\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|$
+Stability in large-scale pretraining: Dozens of approaches address instability during large-model pretraining, including: architectural advances (Pre-LN (Xiong et al., 2020), RMSNorm (Zhang & Sennrich, 2019)), careful initialization (Nguyen & Salazar, 2019; Takase et al., 2023; Nishida et al., 2024), auxiliary loss terms (Max-z loss (Yang et al., 2023)).
+Recent work (OLMo et al., 2024) also explores combining multiple stabilization strategies.
+These measures improve average stability but do not directly prevent abnormal gradients from corrupting optimizer states.
+Gradient/Update Clipping: Gradient and update clipping achieve stability by limiting the magnitude of gradients and parameter updates, preventing excessively large weight updates.
+Global gradient clipping (Pascanu et al., 2013) is prevalent, with innovative approaches like AGC (Brock et al., 2021) and Clippy (Tang et al., 2023), which use model weights to adjust the clipping threshold.
+The SPAM (Huang et al., 2025) method stabilizes the model training process by introducing a momentum reset mechanism and an element-wise gradient clipping strategy based on second-moment estimation.
+Alternatives like Adafactor (Shazeer & Stern, 2018), StableAdamW (Wortsman et al., 2023), and LAMB (You et al., 2019) offer update clipping techniques better suited for stability training of large-scale models.
 Nonetheless, a significant number of loss spikes still occur during the training of large language models, even with the application of these methodologies.
-3 Motivations
-(a) Temporal threshold adaptation necessity: Fixed thresholds (0.4) prevent early spikes but fail later (10k steps), requiring dynamic adjustment.
-(b) Spikes in gradient norms occur at different training steps across parameters, indicating heterogeneous gradient behaviors.
-(c) Per-parameter gradient clipping mitigates loss spikes that global clipping cannot, leading to faster convergence.
-Figure 1: Empirical motivation for AdaGC: (a) Temporal threshold decay necessitates adaptive clipping, (b) Parameter-specific gradient spikes demand localized control, (c) Fine-grained clipping outperforms global approaches.
-In this section, we will introduce some phenomena of gradient clipping that motivate the design of the algorithm.
-Adaptivity: The magnitudes of gradients in neural network training exhibit dynamic variations throughout the learning process.
-This renders fixed gradient clipping thresholds suboptimal for mitigating loss spikes, as demonstrated in Figure 1(a).
-While a threshold of 0.4 successfully mitigates loss spike during the initial training phase of the GPT-2 345M model, it becomes ineffective against the severe loss spike occurring around 10,000 iterations, where a substantially reduced threshold is required for stabilization.
-This phenomenon correlates with the observation that optimal clipping thresholds for convergence speed are strongly influenced by recent gradient norms (see corresponding subfigures).
-Motivated by these findings, we propose an adaptive clipping mechanism using the exponential moving average of gradient magnitudes as the per-step clipping threshold.
-Locality: Figure 1(b) reveals that different parameters exhibit divergent gradient behaviors, with spikes occurring at distinct training phases.
-Using a global clipping threshold can degrade the convergence speed of certain parameters.
-Therefore, we propose replacing global gradient norm clipping with per-parameter norm clipping, where clipping is applied independently to each parameter.
-Furthermore, as illustrated in Figure 1(c), we evaluate the effectiveness of per-parameter clipping thresholds categorized by specific ratios on GPT2-345M.
-The results indicate that, under the same global gradient norm conditions, per-parameter clipping can address spike phenomena that global clipping cannot resolve, and achieve faster convergence speeds.
-These observations collectively motivate our two core design principles: adaptive thresholding to address temporal gradient norm decay, and localized control to handle parameter-wise gradient heterogeneity.
-The dynamic nature of optimal clipping thresholds (Figure 1(a)) necessitates our exponential moving average mechanism for temporal adaptation, while asynchronous parameter gradient spikes (Figure 1(b)) justify per-parameter clipping granularity.
-Together, these principles form the foundation of AdaGC - adaptively adjusting localized thresholds that respect both the temporal evolution and spatial distribution of gradients.
+Due to our gradient-centric perspective, we focus our discussion on clipping-based methods.
+These methods fall into two categories: value-based approaches, which truncate individual gradient components exceeding a predefined limit, and norm-based approaches, which rescale the entire gradient vector only when its overall magnitude exceeds a threshold.
+AdaGC belongs to the norm-based category, leveraging adaptive per-tensor norm thresholds to stabilize training.
+For a comparative summary, see Table 2.
+Figure 2: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the embedding of Llama-2 1.3B during warmup phase.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+3 Motivation: From Root-Cause Diversity to a Unified Gradient-Centric Remedy
+Through a series of experiments (see Figure 1 and Figure 2), we observe that loss spikes encountered under diverse settings consistently coincide with abrupt fluctuations in the gradient norm.
+Comparative analyses further reveal limitations of existing methods such as GlobalGC, AGC, and Clippy: GlobalGC’s static global threshold cannot detect or suppress localized abnormal gradients, allowing outliers to contaminate optimizer states and trigger instability.
+AGC and Clippy focus on controlling parameter updates, leaving internal moments vulnerable to large gradient outliers.
+As discussed in the Introduction (Section  1), loss spikes typically result from a combination of multiple factors.
+While the specific triggers may vary, these loss spikes share a common manifestation: abnormally large gradients are incorporated into the optimizer’s moment estimates, leading to unstable updates.
+Based on these analyses, we propose a unified remedy: regardless of the root cause, instability in large-scale training is best addressed via gradient-centric clipping.
+Specifically, only localized and adaptive clipping, applied before gradients are integrated into the optimizer’s moment estimates, can effectively constrain the influence of outlier gradients.
+We thus distill two key principles for loss spike mitigation: (1) Locality: clip gradients for each parameter tensor individually, avoiding the insensitivity of a global threshold; (2) Adaptivity: dynamically adjust each tensor’s clipping threshold, e.g., using an EMA of its recent gradient norms.
 4 Methodology: AdaGC
 4.1 Preliminaries
-[…]
+Notations.
+Let $x_{t}\in\mathbb{R}^{d}$ denote a parameter vector where $x_{t}^{j}$ represents its $j$-th coordinate for $j\in[d]$.
+We write $\nabla_{x}f(x)$ for the gradient of any differentiable function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and use $u^{2}$ and $u/v$ to denote element-wise square and division operations for vectors $u,v\in\mathbb{R}^{d}$.
+The $\ell_{2}$-norm and $\ell_{\infty}$-norm are denoted by $\|\cdot\|$ and $\|\cdot\|_{\infty}$, respectively.
+For asymptotic comparisons, we write $f=\mathcal{O}(g)$ if $\exists c>0$ such that $f(x)\leq cg(x)$ for all $x$ in the domain.
 Gradient Clipping Fundamentals.
@@
 Standard gradient descent updates follow:
-$\theta_{t}=\theta_{t-1}-\alpha\nabla_{\theta}f(\theta_{t-1},X_{t})$ (1)
-To prevent unstable updates from gradient explosions, gradient clipping (Pascanu et al., 2013) modifies the update rule as:
-$\begin{array}[]{c}\theta_{t}=\theta_{t-1}-\alpha h_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})\\[3.99994pt] \text{ where }h_{t}:=\min\left\{\frac{\lambda_{abs}}{\|\nabla_{\theta}f(\theta_{t-1};X_{t})\|},1.0\right\}\end{array}$ (2)
-Here $\lambda_{\text{abs}}$ is an absolute clipping threshold requiring careful tuning.
-Our work focuses on norm-based clipping (scaling entire gradients exceeding $\lambda_{\text{abs}}$) rather than value-based clipping (element-wise truncation).
+$\small\theta_{t}=\theta_{t-1}-\eta_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})$ (1)
+To prevent unstable updates from gradient explosions, GlobalGC (Pascanu et al., 2013) modifies the update rule as:
+$\small\begin{array}[]{c}\theta_{t}=\theta_{t-1}-\eta_{t}h_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})\\[3.6999pt] \text{ where }h_{t}:=\min\left\{\frac{\lambda_{abs}}{\|\nabla_{\theta}f(\theta_{t-1};X_{t})\|},1.0\right\}\end{array}$ (2)
+Here $\lambda_{abs}$ is an absolute clipping threshold requiring careful tuning, and $\eta_{t}$ is the learning rate.
+Our work focuses on norm-based clipping (scaling entire gradients exceeding $\lambda_{abs}$) rather than value-based clipping (element-wise truncation).
 4.2 Adaptive Gradient Clipping based on Local Gradient Norm
-This section introduces a novel gradient clipping strategy termed Adaptive Gradient Clipping (AdaGC), which distinguishes itself by not relying on a global gradient norm.
-Instead, AdaGC focuses on the local gradient norm of each parameter and utilizes a dynamic adaptive mechanism for gradient clipping.
-The proposed method employs an Exponential Moving Average (EMA) mechanism to maintain smoothed estimates of historical gradient norms per parameter, thus enhancing the accuracy of anomalous gradient detection and enabling independent clipping adjustments tailored to each parameter’s specific conditions.
+This section introduces a novel gradient clipping strategy termed AdaGC, which distinguishes itself by not relying on a global gradient norm.
+Instead, AdaGC focuses on the local gradient norm of each tensor and utilizes a dynamic adaptive mechanism for gradient clipping.
+The proposed method employs an EMA mechanism to maintain smoothed estimates of historical gradient norms per tensor, thus enhancing the accuracy of anomalous gradient detection and enabling independent clipping adjustments tailored to each tensor’s specific conditions.
 EMA is widely used in deep learning, and within AdaGC, it facilitates the balancing of historical and current gradient norms.
 The formulation is as follows:
-$\begin{array}[]{c}\boldsymbol{g}_{t,i}\leftarrow h_{t,i}\cdot\boldsymbol{g}_{t,i},\text{where }h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\},\\[10.00002pt] \gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|.\end{array}$ (3)
-Here, $\lambda_{rel}$ is a predefined relative clipping threshold, $\boldsymbol{g}_{t,i}$ represents the gradient of the $i$-th parameter at time step $t$, and $h_{t,i}$ is a clipping function activated when $\|\boldsymbol{g}_{t,i}\|>\lambda_{rel}\cdot\gamma_{t-1,i}$, thereby scaling the gradient norm to $\lambda_{rel}\cdot\gamma_{t-1,i}$.
+$\small\begin{array}[]{c}\boldsymbol{g}_{t,i}\leftarrow h_{t,i}\cdot\boldsymbol{g}_{t,i},\text{where }h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\},\\[9.24994pt] \gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|.\end{array}$ (3)
+Here, $\lambda_{rel}$ is a predefined relative clipping threshold, $\boldsymbol{g}_{t,i}$ represents the gradient of the $i$-th tensor at time step $t$, and $h_{t,i}$ is a clipping function activated when $\|\boldsymbol{g}_{t,i}\|>\lambda_{rel}\cdot\gamma_{t-1,i}$, thereby scaling the gradient norm to $\lambda_{rel}\cdot\gamma_{t-1,i}$.
 Additionally, $\beta$ is the smoothing coefficient for EMA.
 We consistently incorporate the clipped gradient norm into the historical observations rather than the pre-clipped values.
-Despite its simplicity, AdaGC adaptively adjusts based on the magnitude of each parameter’s gradient norm.
+Despite its simplicity, AdaGC adaptively adjusts based on the magnitude of each tensor’s gradient norm.
 Whenever the gradient norm at a current timestep exceeds a predefined range of average norms within a historical window, it effectively suppresses these outlier gradients.
 However, during the initial stages of model training (e.g., the first 100 steps), the gradient norms are typically large and fluctuate significantly, indicating a substantial decreasing trend.
-Direct application of AdaGC during this period could lead to two issues: first, erroneously accumulating the early large gradient norms into the historical values, resulting in compounded errors; second, compared to traditional global-norm-based methods, AdaGC might delay clipping, thus potentially slowing down the loss reduction.
-To address these issues, we introduce a hyperparameter $T_{start}$(default set to 100), representing a warm-up period during which traditional global gradient norm clipping is applied.
-Additionally, the AdaGC strategy can be seamlessly integrated with various optimizers, such as AdamW (Loshchilov & Hutter, 2017), enhancing its practicality and flexibility.
-Algorithm 1 demonstrates its implementation with the AdamW optimizer.
-Overall, AdaGC offers a more precise and effective gradient management method for large-scale model training, contributing to improved training stability and performance.
+Direct application of AdaGC during this period could lead to two issues: first, erroneously accumulating the early large gradient norms into the historical values, resulting in compounded errors; second, compared to GlobalGC, AdaGC might delay clipping, thus potentially slowing down the loss reduction.
+To address these issues, we introduce a hyperparameter $T_{start}$ (default set to 100), representing a warm-up period during which traditional GlobalGC is applied.
+Additionally, AdaGC is optimizer-agnostic, can be seamlessly integrated with various optimizers, such as AdamW (Loshchilov & Hutter, 2017), Lion (Chen et al., 2024), Muon (Jordan et al., 2024), enhancing its practicality and flexibility.
+Algorithm 1 in Appendix A demonstrates its implementation with the AdamW optimizer.
+Table 3: Zero-shot accuracy of AdaGC on Llama-2 7B under different hyperparameters.
+$\lambda_{rel}$ $\beta$ 0.98 0.985 0.99 0.999
+1.03 50.06 50.92 50.95 50.96
+1.04 48.88 50.59 51.04 50.76
+1.05 51.01 49.95 50.57 50.74
+Table 4: Two-shot accuracy of AdaGC on Llama-2 7B under different hyperparameters.
+$\lambda_{rel}$ $\beta$ 0.98 0.985 0.99 0.999
+1.03 52.31 52.68 53.13 53.42
+1.04 52.68 53.01 53.47 52.85
+1.05 52.68 52.67 51.96 53.51
+4.3 Memory, Computation, and Communication
+Memory.
+As a tensor-wise method, AdaGC maintains an EMA of gradient norms for each parameter tensor, requiring storage of a single 32-bit float (4 bytes) per tensor.
+For ERNIE models, the total additional memory overhead has complexity of $\mathcal{O}((9+3E)\times L+3)$, where $L$ and $E$ denote the number of transformer layers and experts, respectively.
+Specifically, this includes four tensors from the attention module per layer, $3\times(1+E)$ tensors from the shared and router experts per layer, and two RMSNorm tensors per layer; plus one tensor each for the embedding layer, the final layer normalization, and the language modeling head.
+In practice, this added memory footprint is negligible compared to the overall memory requirements of large-scale model training.
+Computation.
+The computational cost of computing $\ell_{2}$ norms is the same for both AdaGC and GlobalGC.
+The difference is that GlobalGC applies a uniform scaling to all gradients, while AdaGC scales each gradient tensor independently.
+Communication.
+In setups involving data parallelism (DP), tensor parallelism (TP), and pipeline parallelism (PP), GlobalGC requires an all-reduce operation across all DP, TP, and PP groups to aggregate the global norm.
+In contrast, AdaGC only needs an all-reduce within each TP group to compute per-tensor local norms.
+This design substantially reduces communication overhead, offering increasing benefits as model and cluster sizes grow.
+5 Experiments
+5.1 Experimental Setup
+[…]
+Comparison Methods.
+We focus on clipping-based methods and compare gradient and update clipping baselines, including GlobalGC (Pascanu et al., 2013), Gradient Value Clipping (ClipByValue), AGC (Brock et al., 2021), and Clippy (Tang et al., 2023).
+We also evaluate recent methods, including SPAM (Huang et al., 2025), Scaled Embed (Takase et al., 2023), and WeSaR (Nishida et al., 2024).
+Results are in Appendix D.2 Table 12.
+Training Details.
+Pre-training large-scale models is typically resource-intensive.
+In this study, our primary focus was to investigate training instability rather than to pursue ultimate accuracy.
+To facilitate multiple experimental runs, we conducted 9,000 training steps over 36 billion tokens for both Llama-2 1.3B and 7B, 36,000 steps over 36 billion tokens for Mixtral 8x1B, and 60,000 steps over 1 trillion tokens for ERNIE 10B-A1.4B, the latter additionally serving to validate the long-term stability of AdaGC.
+Further details regarding hyperparameters can be found in Table 9 in Appendix B.
+Evaluation Metrics.
+To quantitatively assess training stability, we follow (OLMo et al., 2024; Karpathy, 2024) and adopt the spike score as an objective metric.
+Specifically, the spike score is defined as the percentage of values in a time series that deviate by at least ten standard deviations from a rolling average of the preceding 1,000 values.
+This metric is primarily applied to training loss to detect sudden instabilities.
+Additionally, we evaluate performance using the training loss and validation perplexity (PPL) curves, as well as standard benchmark results, to provide a comprehensive assessment of convergence efficiency and model quality.
+Standard Benchmark.
+We conducted a comprehensive evaluation of the model’s zero-shot and two-shot capabilities across seven well-established benchmarks: ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), and MMLU (Hendrycks et al., 2020).
+Following standard practice (Zhang et al., 2025), we report accuracy norm for ARC-E, ARC-C, HellaSwag, OBQA, and SciQ, as well as standard accuracy for all other tasks.
+For ERNIE 10B-A1.4B, which has been trained on 1T tokens, we evaluate its general abilities on a range of benchmarks, including MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2021), and HumanEval (Chen et al., 2021).
+These benchmarks assess the model’s enhanced capabilities in performing diverse downstream tasks, such as examination, reasoning, factuality, and coding.
+(a) Llama-2 7B training dynamics.
+(b) Mixtral 8x1B training dynamics.
+Figure 3: Large language model training analysis: Llama-2 7B and Mixtral 8x1B model comparison shows AdaGC’s loss spike elimination and performance gains.
+5.2 Critical Hyperparameter Selection
+We systematically evaluated two key hyperparameters in AdaGC: the EMA coefficient $\beta$ and the relative clipping threshold $\lambda_{rel}$.
+Specifically, we performed a grid search on the Llama-2 7B model to optimize these two hyperparameters, using zero-shot and two-shot performance across multiple tasks as evaluation metrics.
+As shown in Tables 4 and 4, the best performance was achieved when $\lambda_{rel}=1.04$ and $\beta=0.99$.
+We therefore adopted this configuration as the default setting for subsequent experiments and terminated further hyperparameter search.
+In addition, as observed in Tables 4 and 4, AdaGC’s performance remains relatively stable across different hyperparameter values, suggesting that the method is robust to hyperparameter variations.
+5.3 Main Experimental Results
+[…]
+Table 5: The Zero-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks.
+Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
+Llama-2 1.3B GlobalGC 43.18 25.68 57.19 46.62 30.20 69.97 52.64 22.97 68.40 46.32
+ClipByValue 42.17 25.68 59.94 44.11 30.40 69.59 53.28 22.99 68.00 46.24
+Clippy 41.71 24.66 56.51 45.43 30.00 69.21 54.85 22.90 67.50 45.86
+AdaGC 42.09 25.51 58.01 47.29 30.40 69.70 52.33 22.98 68.70 46.33
+Llama-2 7B GlobalGC 49.49 27.56 56.30 56.06 33.60 74.59 55.33 23.12 71.20 49.69
+ClipByValue 46.21 26.88 57.03 53.49 33.20 71.65 53.59 23.36 70.50 48.43
+AGC 48.15 28.16 52.87 55.47 32.80 72.74 57.85 24.33 71.70 49.34
+Clippy 47.69 27.73 57.46 53.34 32.40 72.74 54.38 25.36 73.40 49.39
+AdaGC 49.58 28.92 57.28 57.94 32.80 74.32 58.09 23.62 76.60 51.01
+Mixtral 8x1B GlobalGC 44.70 25.94 56.57 53.08 33.00 71.60 54.70 22.91 67.20 47.74
+AdaGC 46.68 26.37 58.93 55.85 32.20 73.12 54.38 23.22 70.30 49.01
+Table 6: Evaluation results of ERNIE 10B-A1.4B on multiple benchmarks after 21,000 (350B tokens) and 60,000 (1T tokens) training steps, comparing different optimization configurations.
+Steps (tokens) Method AdamW eps MMLU GSM8K BBH TruthfulQA HumanEval Avg.
+21k (350B) GlobalGC 1e-8 47.75 28.35 28.80 22.02 19.51 28.09
+GlobalGC 1e-15 39.11 21.46 29.35 23.39 15.24 25.71
+AdaGC 1e-15 42.07 25.32 27.89 24.92 20.73 28.19
+60k (1T) GlobalGC 1e-8 48.61 39.88 30.84 30.73 22.56 34.52
+GlobalGC 1e-15 48.48 40.79 30.59 28.29 23.78 34.38
+AdaGC 1e-15 48.70 36.01 31.38 35.02 22.56 34.73
+[…]
+Table 7: Comparison of spike scores for various models under different clipping methods.
+Model Llama-2 7B Mixtral 8x1B ERNIE 10B-A1.4B
+Method GlobalGC ClipByValue AGC Clippy AdaGC GlobalGC AdaGC GlobalGC AdaGC
+Total Steps 9K 9K 9K 9K 9K 36K 36K 21K 21K
+Num Spikes 3 9 8 3 0 52 0 2 0
+Spike Score (%) 0.0333 0.1000 0.0889 0.0333 0.0000 0.0144 0.0000 0.0100 0.0000
+Table 8: GPU hours under the same configuration. DPS denotes distributed parallel strategies.
+Model Llama-2 1.3B Llama-2 7B Mixtral 8x1B ERNIE 10B-A1.4B
+DPS DP=256, TP=1, PP=1 DP=32, TP=2, PP=1 DP=256, TP=1, PP=1, EP=1 DP=64, TP=1, PP=4, EP=8
+Steps 9K 9K 36K 21K
+GlobalGC 513.0 1468.2 2060.8 22922
+AdaGC 511.6 1402.4 2035.2 22572
+[…]
+5.4 Optimizer Compatibility: Muon and Lion
+AdaGC is an optimizer-agnostic gradient clipping method that can be seamlessly integrated not only with AdamW, but also with other optimizers.
+To verify the generality of AdaGC, we conducted experiments on both LLM and VLM tasks by combining Llama-2 1.3B and CLIP ViT-Base models with the Muon and Lion optimizers, respectively, and compared them against GlobalGC.
+Although no loss spikes were observed under either experimental setting, AdaGC consistently demonstrated strong compatibility and generalization.
+In downstream zero-shot average accuracy, AdaGC outperformed GlobalGC by 0.14% (47.18% vs.
+47.04%) with Muon and by 0.16% (40.81% vs.
+40.65%) with Lion.
+These results further confirm that AdaGC can be effectively applied across different optimizers, providing stable training and improved downstream performance.
+5.5 End-to-End training Wall-clock
+Table 8 compares the GPU hours required for training various models using different distributed parallelism strategies.
+Compared to GlobalGC, AdaGC reduces end-to-end GPU hours by 0.27% on Llama-2 1.3B, 4.48% on Llama-2 7B, 1.24% on Mixtral 8x1B, and 1.53% on ERNIE 10B-A1.4B, thanks to reduced communication overhead.
+This highlights AdaGC’s additional communication and efficiency benefits in large-scale distributed training.
+5.6 Ablation Study
+We conduct systematic ablation studies across three critical dimensions of AdaGC: (1) EMA gradient norm initialization strategies, (2) GlobalGC warm-up steps, (3) adaptivity efficacy, and (4) locality granularity.
+(a) Training dynamics.
+(b) Average accuracy.
+Figure 4: AdaGC with Muon on Llama-2 1.3B.
+(c) Training dynamics.
+(d) Average accuracy.
+Figure 5: AdaGC with Lion on CLIP ViT-Base.
+(a) $\gamma_{t,i}$ initialization.
+(b) $T_{start}$ warm-up.
+(c) Adaptivity, locality.
+Figure 6: Training dynamics of AdaGC ablations, showing (a) the influence of different EMA initialization strategies; (b) the impact of the GlobalGC warm-up steps $T_{start}$; and (c) the effects of adaptivity and locality granularity on gradient clipping efficacy and final loss.
+EMA Initialization Strategy.
+The initialization of EMA gradient norms requires careful design due to large initial gradient fluctuations during early training phases (first 100 steps).
+We evaluate five initialization variants: The default AdaGC strategy employs GlobalGC during warm-up while tracking minimum per-parameter norms ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$).
+Comparative approaches include: (1) norm initialization without GlobalGC warm-up (directly using $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$ from step 0), (2) constant initialization ($\gamma_{0,i}\in\{0.5,1.0\}$), and (3) thresholded initialization ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,0.1)$).
+Figure 6(a) demonstrates that while all variants eliminate loss spikes, convergence quality varies within 0.36%.
+The default strategy achieves optimal final loss (2.9708 vs 2.9725 for next-best), showing that GlobalGC-guided warm-up better preserves parameter update consistency than direct initialization.
+This establishes the importance of phased initialization for gradient norm adaptation.
+Warm-up Steps $T_{start}$.
+To further investigate whether the choice of GlobalGC warm-up steps $T_{start}$ has a significant impact and to provide practical guidance for practitioners, we additionally evaluate $T_{start}=\{0,50,100,150,200,500,1000,2000\}$.
+The results in Figure 6(b) show that $T_{start}=100$ consistently achieves the best performance.
+According to the EMA initialization formula $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$, an excessively large $T_{start}$ accumulates lower $\gamma_{t,i}$ values due to early training dynamics, which may lead to over-clipping and suppressed convergence in later training.
+Conversely, an overly small $T_{start}$ accumulates larger $\gamma_{t,i}$ values, which may delay clipping and hinder timely suppression of abnormal gradients.
+In contrast, $T_{start}=100$ introduces negligible additional overhead for large-scale training while providing consistently stable performance improvements.
+Adaptivity Efficacy and Locality Granularity.
+We conduct three sets of ablation experiments to evaluate the adaptivity and locality of AdaGC.
+The baseline uses GlobalGC (no adaptivity, no locality) with a fixed threshold of 1.0.
+In comparison, we examine (1) adaptive global gradient norm clipping (Global AdaGC, adaptive but non-local), which employs a single adaptive threshold for the entire model, (2) tensor-wise gradient norm clipping (TensorWiseGC, local but non-adaptive), which allocated each tensor’s fixed clipping threshold proportionally to its parameter count relative to the entire model, and (3) tensor-wise adaptation (AdaGC, adaptive and local), which adjusts thresholds independently for each tensor.
+As shown in Figure 6(c), Global AdaGC reduces but does not completely eliminate spike events (1 event vs.
+0 for tensor-wise) and yields a 0.25% higher final loss (2.9639 vs.
+2.9712).
+Although TensorWiseGC also mitigates loss spikes, it noticeably slows down convergence and requires careful per-tensor threshold tuning to perform well.
+These results demonstrate that tensor-wise adaptive clipping provides both greater spike suppression and lower loss than other approaches.
+6 Conclusion
+The factors triggering loss spikes in large-scale pretraining are diverse and remain an open research problem, with no unified solution to date.
+Unlike prior work that seeks to identify root causes, we focus on a gradient-centric remedy and introduce AdaGC, an adaptive per-tensor gradient clipping method that prevents abnormal gradients from contaminating optimizer states.
+This approach ensures smoother updates and effectively eliminates loss spikes.
+Extensive experiments demonstrate that AdaGC delivers robust and stable training across both dense and MoE models, from 1.3B to 10B parameters, consistently reducing spike scores to zero and improving benchmark performance.
+Our results highlight AdaGC as a simple and effective solution for stable large-scale model pretraining.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix Contents
+(Run LaTeX twice to generate Appendix Contents.)
+Appendix A Pseudocode of AdamW with AdaGC
+Algorithm 1 presents the pseudocode of AdamW integrated with AdaGC.
+For clearer exposition, we highlight different components according to their origins: orange indicates the procedures inherited from the original GlobalGC algorithm, while blue is used to denote the new contributions and modifications introduced by AdaGC.
+Specifically, the GlobalGC steps include the global gradient clipping implemented via the scaling factor and the use of the clipped gradient in subsequent moments.
+The AdaGC components mainly comprise adaptive per-parameter clipping, the initialization and update of the adaptive threshold $\gamma_{t,i}$, and the warm-up strategy governed by $T_{start}$.
 Algorithm 1 AdamW with AdaGC
-1:  given: $\{\alpha_{t}\}_{t=1}^{T},\lambda_{w},\epsilon_{1},\beta_{1},\beta_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta}\in[0,1),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\lambda_{abs}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{start}}$
+1:  given: $\{\eta_{t}\}_{t=1}^{T},\lambda_{w},\epsilon,\beta_{1},\beta_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta\in[0,1)},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\lambda_{abs}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{start}}$
 2:  initialize: $\boldsymbol{\theta}_{0},m_{0}\leftarrow 0,v_{0}\leftarrow 0,t\leftarrow 0$
@@
 5:  if $t<T_{start}$ then
-6:       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h_{t}=\min\left\{\frac{\lambda_{abs}}{\|\boldsymbol{g}_{t}\|},1.0\right\}}$
-7:       ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}=h_{t}\cdot\boldsymbol{g}_{t}}$
+6:   $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}h_{t}=\min\left\{\frac{\lambda_{abs}}{\|\boldsymbol{g}_{t}\|},1.0\right\}$}
+7:   $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}=h_{t}\cdot\boldsymbol{g}_{t}$}
 8:   for $i\in|\boldsymbol{\theta}|$ do
-9:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\min\left\{\gamma_{t-1,i},\|\widehat{\boldsymbol{g}_{t,i}}\|\right\},\gamma_{0,i}=\|\widehat{\boldsymbol{g}_{1,i}}\|}$
+9:    $^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\min\left\{\gamma_{t-1,i},\|\widehat{\boldsymbol{g}_{t,i}}\|\right\},\gamma_{0,i}=\|\widehat{\boldsymbol{g}_{1,i}}\|}$
 10:   end for
@@
 14:    $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t,i}}=h_{t,i}\cdot\boldsymbol{g}_{t,i}$}
-15:          ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\widehat{\boldsymbol{g}_{t,i}}\|}$
+15:    $^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\widehat{\boldsymbol{g}_{t,i}}\|}$
 16:   end for
 17:  end if
-18:    $\boldsymbol{m}_{t}=\beta_{1}\boldsymbol{m}_{t-1}+(1-\beta_{1}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}}$
-19:    $\boldsymbol{v}_{t}=\beta_{2}\boldsymbol{v}_{t-1}+(1-\beta_{2}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t}}}^{2}$
+18:  $\boldsymbol{m}_{t}=\beta_{1}\boldsymbol{m}_{t-1}+(1-\beta_{1}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}}$
+19:  $\boldsymbol{v}_{t}=\beta_{2}\boldsymbol{v}_{t-1}+(1-\beta_{2}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}^{2}}$
 20:  $\widehat{\boldsymbol{m}_{t}}=\boldsymbol{m}_{t}/(1-\beta_{1}^{t}),\quad\widehat{\boldsymbol{v}_{t}}=\boldsymbol{v}_{t}/(1-\beta_{2}^{t})$
-21:    $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\alpha_{t}\lambda_{w}\boldsymbol{\theta}_{t-1}-\alpha_{t}\widehat{\boldsymbol{m}_{t}}/(\sqrt{\widehat{\boldsymbol{v}_{t}}}+\epsilon_{1})$
+21:  $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta_{t}\lambda_{w}\boldsymbol{\theta}_{t-1}-\eta_{t}\widehat{\boldsymbol{m}_{t}}/(\sqrt{\widehat{\boldsymbol{v}_{t}}}+\epsilon)$
 22:  until $\boldsymbol{\theta}_{t}$ not converge
-4.3 Convergence Analysis
-In this section, we give the convergence guarantee of Adam with AdaGC stated as follows:
-Theorem 4.1.
+Appendix B Hyper-Parameters
+B.1 Model Hyper-Parameters
+Table 9 summarizes the model hyper-parameters used for all experiments.
+For each model, we report the core architecture settings (such as number of layers, hidden dimension, attention heads, and feedforward dimension), MoE-related configurations, and main optimization hyper-parameters (including learning rate, warmup, weight decay, and Adam parameters).
+Clipping thresholds $\lambda_{abs}$, $\lambda_{rel}$, and momentum $\beta$ are also listed, in correspondence with the techniques discussed in the main text.
+All experiments use a batch size and sequence length as shown, and we employ bfloat16 precision for most models except ERNIE, which uses float8.
+The symbol ‘–’ indicates settings not applicable to a specific architecture.
+Table 9: Hyper-parameters used in our LLMs experiments.
+$\lambda_{abs}$ represents the absolute global clipping threshold of GlobalGC.
+$\lambda_{rel}$ and $\beta$ represent the relative clipping threshold and the momentum of our AdaGC, respectively.
+The symbol ‘–’ indicates that the parameter is not applicable.
+Model Llama-2 1.3B Llama-2 7B ERNIE 10B-A1.4B Mixtral 8x1B
+Precision bfloat16 bfloat16 float8 bfloat16
+Num layers 24 32 25 24
+Hidden dim size 2048 4096 2560 2048
+FFN dim size 5461 11008 1024 5632
+Num attention heads 32 32 20 32
+Num key value heads 32 32 4 4
+Attention bias
+Num shared experts - - 1 0
+Num router experts - - 48 8
+Num experts per token - - 3 2
+Sequence length 2048 2048 4096 2048
+Batch size 2048 2048 4096 512
+Iterations 9000 9000 21000 36000
+Learning rate $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$
+LR decay cosine cosine wsd cosine
+Warmup iterations 2000 2000 2000 500
+Weight decay 0.1 0.1 0.1 0.1
+Adam $\beta_{1}$ 0.90 0.90 0.90 0.90
+Adam $\beta_{2}$ 0.95 0.95 0.95 0.999
+$\lambda_{abs}$ 1.0 1.0 1.0 1.0
+$\lambda_{rel}$ 1.04 1.04 1.04 1.04
+$\beta$ 0.99 0.99 0.99 0.99
+B.2 Clipping Hyper-Parameters
+For other clipping methods, we primarily followed the recommended default settings from prior work, and performed limited tuning only when necessary to ensure a fair comparison.
+Specifically:
+  • GlobalGC: We used the commonly adopted global clipping threshold $\lambda_{abs}=1.0$ in large-scale pretraining.
+  • ClipByValue: Following the SPAM (Huang et al., 2025) setting, we set the clipping threshold to $\lambda_{abs}=1e-3$.
+  • AGC: We performed small-range tuning over $\lambda_{rel}\in\{1e-2,1e-3,1e-4\}$ to find the best setting.
+  • Clippy: We tuned over $\lambda_{abs}\in\{0.1,0.3,0.5\}$ and $\lambda_{rel}\in\{1e-2,1e-3,1e-4\}$ to select the optimal combination.
+  • SPAM: We adopted the default hyperparameters recommended for standard pretraining in the original paper, which were reported to perform well across diverse settings.
+Specifically, we set the interval to $\Delta T=500$, the threshold to $\theta=5000$, and the warmup steps to $N=150$.
+The final hyper-parameters used for other clipping methods are summarized in Table 10.
+Table 10: Hyper-parameters for other clipping methods.
+Method Hyperparameters
+GlobalGC $\lambda_{abs}=1.0$
+ClipByValue $\lambda_{abs}=1e-3$
+AGC $\lambda_{rel}=1e-3$
+Clippy $\lambda_{abs}=0.5,\lambda_{rel}=1e-2$
+SPAM $\Delta T=500$, $\theta=5000$, $N=150$
+Appendix C Experimental Details for CLIP
+To further investigate the optimizer compatibility of AdaGC, we evaluated its effect on large-scale vision-language model pre-training, focusing on the CLIP ViT-Base model (Radford et al., 2021) with the Lion optimizer (Chen et al., 2024).
+The model comprises 151 million parameters and is trained on the LAION-400M (Schuhmann et al., 2021) dataset.
+Training is conducted for 20,000 steps, covering 320M image-text pairs.
+The key training hyper-parameters are as follows: a learning rate of 0.002, weight decay of 0.2, and batch size of 32,768.
+We employ patch-dropout with a drop rate of 0.5 (Li et al., 2023), following recent best practices (Wortsman et al., 2023).
+The learning rate is linearly warmed up for the first 5,000 steps (Goyal et al., 2017), and subsequently decayed according to a cosine schedule (Loshchilov & Hutter, 2016).
+Following pre-training, we report downstream zero-shot evaluation results on the ImageNet (Russakovsky et al., 2015) validation set. The results are shown in Figure 5 in the main text.
+Appendix D More Evaluation Results
+D.1 Results on Downstream Benchmarks
+The Two-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks are presented in Table 11.
+Table 11: The Two-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks. The best scores in each column are bolded. HellaSw. = HellaSwag, W.G. = WinoGrande.
+Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
+Llama-2 1.3B GlobalGC 47.26 25.60 50.31 46.44 32.20 69.64 52.33 25.07 77.80 47.41
+ClipByValue 47.10 25.77 56.54 43.97 30.00 68.88 52.96 26.09 77.20 47.61
+Clippy 46.55 25.85 49.76 45.71 30.00 70.02 53.20 25.69 77.70 47.16
+AdaGC 46.04 26.19 49.72 47.51 31.00 69.70 54.38 24.98 78.50 47.56
+Llama-2 7B GlobalGC 55.81 28.58 60.70 56.54 33.00 73.72 56.75 25.51 83.20 52.64
+ClipByValue 51.94 26.88 57.55 53.36 32.40 72.31 54.14 26.63 81.60 50.75
+AGC 52.95 28.67 56.15 55.69 35.40 73.07 56.43 26.88 82.80 52.00
+Clippy 52.86 29.10 56.48 53.76 31.80 73.07 55.72 26.03 82.60 51.27
+AdaGC 56.86 29.61 59.36 57.89 33.60 73.99 57.62 26.46 85.90 53.47
+Mixtral 8x1B GlobalGC 50.34 27.39 58.81 52.96 34.20 71.16 54.06 25.37 79.90 50.47
+AdaGC 53.83 28.42 58.69 55.66 33.80 73.07 54.14 25.12 81.80 51.61
+D.2 Results of Other Baseline Methods
+Table 12: The Zero-Shot evaluation results of Llama-2 1.3B/7B models on standard benchmarks.
+Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
+acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
+Llama-2 1.3B WeSaR-GlobalGC 43.56 25.17 59.94 45.08 30.00 70.29 52.96 22.90 65.80 46.19
+SPAM 42.05 24.83 59.60 42.82 30.00 69.31 52.17 23.02 66.40 45.58
+ScaledEmbed-GlobalGC 42.21 25.51 59.66 45.50 31.80 70.02 53.28 23.22 65.20 46.27
+AdaGC 42.09 25.51 58.01 47.29 30.40 69.70 52.33 22.98 68.70 46.33
+Llama-2 7B WeSaR-GlobalGC 49.75 27.22 56.12 55.38 33.80 73.39 56.27 23.02 71.40 49.59
+SPAM 48.53 25.77 60.34 51.89 32.60 72.03 54.54 22.95 71.00 48.85
+ScaledEmbed-GlobalGC 48.57 26.71 60.89 54.32 32.60 72.25 55.33 23.66 70.50 49.42
+AdaGC 49.58 28.92 57.28 57.94 32.80 74.32 58.09 23.62 76.60 51.01
+Table 13: Comparison of spike scores for various models under different methods.
+Model Llama-2 1.3B Llama-2 7B
+Method WeSaR-GlobalGC SPAM ScaledEmbed-GlobalGC AdaGC WeSaR-GlobalGC SPAM ScaledEmbed-GlobalGC AdaGC
+Total Steps 9K 9K 9K 9k 9K 9K 9K 9K
+Num Spikes 2 0 10 0 1 3 8 0
+Spike Score (%) 0.0222 0.0000 0.1111 0.0000 0.0111 0.0333 0.0889 0.0000
+In addition to the clipping-based baselines discussed in the main text, we also compare AdaGC with several recent methods that aim to improve the stability and generalization of large language model (LLM) training, including SPAM (Huang et al., 2025), Scaled Embed (Takase et al., 2023), and WeSaR (Nishida et al., 2024).
+The detailed results under the zero-shot setting and spike score are summarized in Table 12 and 13.
+The training dynamics are shown in Figures 8 and 9.
+Among these methods, SPAM is designed to stabilize training by adjusting the optimizer’s behavior, while Scaled Embed and WeSaR focus on initialization or embedding scaling strategies to suppress loss spikes.
+Our experiments show that, although some of these methods can partly mitigate instability or improve certain metrics, AdaGC generally achieves higher stability and better final performance across model scales.
+Notably, while WeSaR is also effective at suppressing loss spikes, its reliance on special parameter initialization limits its applicability to from-scratch training.
+In contrast, AdaGC works reliably under both from-scratch and resumed training regimes, providing stronger flexibility.
+Overall, these results demonstrate AdaGC’s superior robustness and generalization compared to other non-clipping baselines.
+Appendix E More Visualization Results
+E.1 Training Dynamics
+Figure 7: Llama-2 1.3B training dynamics of clipping methods.
+Figure 8: Llama-2 1.3B training dynamics of other baseline methods.
+Figure 9: Llama-2 7B training dynamics of other baseline methods.
+E.2 Optimizer State Dynamics
+Figure 10: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the embedding of Llama-2 1.3B.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+Figure 11: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_3_self_attention_query_key_value of Llama-2 1.3B.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+Figure 12: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the LMHead of Llama-2 1.3B.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+Figure 13: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_final_layernorm of Llama-2 1.3B.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+Figure 14: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_0_self_attention_query_key_value of Llama-2 1.3B.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+Figure 15: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_23_input_layernorm of Llama-2 1.3B.
+Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC.
+The black curve in each plot shows the loss trajectory.
+Appendix F Convergence Analysis and Proof
+Any operation that modifies gradients may potentially result in non-convergence.
+In this section, rather than providing a theoretical guarantee that AdaGC eliminates loss spikes, we present the convergence guarantee for Adam with AdaGC, stated as follows:
+Theorem F.1.
 Under mild assumptions, by selecting $\alpha_{t}=\mathcal{O}(1/\sqrt{T})$, $\beta_{2}=1-\mathcal{O}(1/T)$ and $\beta_{1}<\sqrt{\beta_{2}}$, when $\tau$ is randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities, it holds that
 $\mathbb{E}\|\nabla f(\theta_{\tau})\|^{2}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
-Theorem 4.1 shows that even with local clipped gradient, Adam with AdaGC can converge in the same rate as vanilla Adam.
-Due to the limited space the formal assumptions, theorem statement with detailed proof can be found in Appendix A.
-5 Experiments
-This paper primarily focuses on the pre-training of large language models, using Llama-2 (Touvron et al., 2023b) as our primary experimental subject.
-Experiments were conducted using NVIDIA GPU A100 40G.
-Moreover, our proposed method demonstrates strong generalization across various VLM models, achieving robust results.
-Table 1: Zero-shot performance of Llama-2 7B on WikiText and LAMBADA datasets.
-Metrics GlobalGC AdaptiveGradientClip StableAdamW AdaGC-Shard AdaGC
-WikiText Word PPL $\downarrow$ 21.173 21.467 21.733 20.878 20.434
-Byte PPL $\downarrow$ 1.770 1.774 1.778 1.765 1.758
-LAMBADA ACC $\uparrow$ 49.35 ($\pm$ 0.70) 49.27 ($\pm$ 0.70) 48.96 ($\pm$ 0.70) 47.76 ($\pm$ 0.70) 49.49 ($\pm$ 0.70)
-PPL $\downarrow$ 11.111 ($\pm$ 0.334) 11.509 ($\pm$ 0.354) 11.774 ($\pm$ 0.363) 11.518 ($\pm$ 0.341) 10.515 ($\pm$ 0.315)
-5.1 Experimental Setup
-Models and Datasets: We evaluated the efficacy of AdaGC on various Llama-2 models including Tiny, 7B and 13B.
-The C4-en (Raffel et al., 2020) dataset, a clean corpus of English text extracted from Common Crawl, served as our pre-training dataset.
-We assessed model performance by computing perplexity on the WikiText (Merity et al., 2016) and LAMBADA (Paperno et al., 2016) datasets.
-We also conducted experiments on GPT-2 (Radford et al., 2019) and CLIP (Radford et al., 2021) models to further validate AdaGC’s generalization capabilities.
-Comparison Methods: We selected several techniques including Global Gradient Clipping (GlobalGC for short) (Pascanu et al., 2013), Adaptive Gradient Clipping (Brock et al., 2021), StableAdamW (Wortsman et al., 2023a), and Scaled Embed (Takase et al., 2023) from different approaches such as gradient clipping, update clipping, and parameter initialization for comparison.
-Training Details: Pre-training large scale models, particularly Llama-2 7B and 13B, is typically resource-intensive.
-Like StableAdamW (Wortsman et al., 2023a) and Scaled Embed (Takase et al., 2023), our primary focus was to explore training instability rather than achieve ultimate accuracy.
-For ease of multiple experiments, we conducted 9,000 training steps on 36 billion tokens for both Llama-2 7B and 13B models, and 36,000 steps on 36 billion tokens for the Llama-2 Tiny model.
-For additional details on the hyperparameters, please refer to the Appendix B.
-Critical Hyperparameter Selection.
-We conducted a systematic study to evaluate the two critical hyperparameters in AdaGC: the EMA coefficient $\beta$ and the relative clipping threshold $\lambda_{rel}$.
-Our analysis utilized coordinate descent (Bertsekas, 1997) for hyperparameter optimization on the Llama-2 Tiny model.
-Initially, we examined the relative clipping threshold $\lambda_{rel}$ with a fixed EMA coefficient $\beta=0.98$, testing the values $\{1.01,1.05,1.10\}$.
-As depicted in Figure 2(a), $\lambda_{rel}=1.05$ achieved optimal performance by effectively balancing gradient regulation effectiveness (compromised at $\lambda_{rel}\geq 1.10$) and update stability (degraded at $\lambda_{rel}\leq 1.01$).
-With this optimal threshold established, we proceeded to investigate the EMA smoothing factor $\beta\in\{0.95,0.98,0.99\}$.
-Figure 2(b) demonstrates that $\beta=0.98$ provides superior adaptation speed while maintaining historical consistency, yielding the best performance.
-Consequently, we adopted the parameter settings $\lambda_{rel}=1.05$ and $\beta=0.98$ for all subsequent experiments reported in this paper.
-(a) Relative clipping threshold $\lambda_{\text{rel}}$.
-(b) EMA coefficient $\beta$.
-Figure 2: Study of AdaGC’s hyperparameters on Llama-2 Tiny.
-(a) Relative clipping threshold analysis demonstrates $\lambda_{\text{rel}}=1.05$ achieves optimal gradient regulation.
-(b) EMA coefficient analysis reveals $\beta=0.98$ best balances historical consistency with rapid adaptation.
-5.2 Experimental Results
-(a) Llama-2 7B training dynamics: AdaGC (green) eliminates loss spikes observed in baseline methods (GlobalGC-blue, ScaledEmbed-orange, AdaptiveGradientClip-purple, StableAdamW-cyan), achieving final training loss 0.02 lower than GlobalGC.
-Validation perplexity (right) shows correlated improvement with stable gradient norms (middle).
-(b) Llama-2 13B scalability analysis: AdaGC maintains superior stability at scale, reducing training loss and validation perplexity by 0.64% and 1.47% respectively versus GlobalGC, demonstrating method scalability.
-Figure 3: Large language model training analysis: (a) 7B model comparison shows AdaGC’s loss spike elimination and performance gains, (b) 13B results demonstrate method scalability.
-Our comprehensive evaluation demonstrates AdaGC’s effectiveness across model scales and architectures.
-Figure 3 presents comparative results on Llama-2 7B and 13B models, analyzing training dynamics through loss trajectories, gradient norms, and validation perplexities.
-For the 7B model, baseline methods (GlobalGC, Scaled Embed, Adaptive Gradient Clipping, and StableAdamW) exhibit frequent loss spikes during training, while AdaGC completely eliminates these instability events.
-This stability translates to measurable performance gains, with AdaGC achieving a 0.02 lower final training loss (2.28 vs 2.30) and 2.1% reduced validation perplexity compared to the strongest baseline (GlobalGC).
-The approach demonstrates strong scalability on the Llama-2 13B model, where AdaGC eliminates loss spikes entirely compared to GlobalGC.
-Qualitative analysis reveals AdaGC’s superior outlier gradient management, maintaining smoother loss trajectories throughout training.
-This enhanced stability yields concrete performance improvements: 0.65% lower training loss (absolute difference of 0.0146) and 1.47% reduced validation perplexity relative to GlobalGC.
-Downstream zero-shot evaluation on WikiText and LAMBADA datasets (Table 1) confirms the practical benefits of stable training.
-AdaGC achieves state-of-the-art performance across all benchmarks, reducing WikiText perplexity by 3.5% (20.434 vs 21.173) while improving LAMBADA accuracy by 0.14 percentage points (49.49% vs 49.35%) compared to GlobalGC.
-These results establish a direct correlation between training stability and final model quality.
-To validate architectural generality, we extend evaluation to vision-language pretraining with CLIP ViT-Base.
-As shown in Figure 4(a), AdaGC completely eliminates loss spikes observed in vision-specific stabilization methods while achieving 25% faster convergence than StableAdamW (reaching target loss in 15k vs 20k steps).
-The improved training dynamics translate to superior zero-shot recognition performance (Figure 4(b)), with AdaGC demonstrating consistent accuracy gains across modalities - including a 0.27 percentage point improvement on ImageNet-1K (39.84% vs 39.57%) over GlobalGC.
-Figure 4: CLIP ViT-Base results: (a) Training loss trajectory comparison.
-AdaGC achieves 25% faster convergence (15k vs 20k steps) with complete loss spike elimination compared to StableAdamW.
-(b) Zero-shot recognition performance.
-Final accuracy improves 0.27pp (39.84% vs 39.57%) on ImageNet-1K, demonstrating cross-modal generalization.
-5.3 Ablation Study
-We conduct systematic ablation studies across three critical dimensions of AdaGC: (1) EMA gradient norm initialization strategies, (2) adaptivity efficacy, and (3) locality granularity.
-The adaptivity property has been empirically validated through experimental analysis (Section 3) and comprehensive experimental results across all benchmarks.
-EMA Initialization Strategy.
-The initialization of EMA gradient norms requires careful design due to large initial gradient fluctuations during early training phases (first 100 steps).
-We evaluate five initialization variants on GPT-2 345M: The default AdaGC strategy employs GlobalGC during warm-up while tracking minimum per-parameter norms ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$).
-Comparative approaches include: (1) norm initialization without GlobalGC warm-up (directly using $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$ from step 0), (2) constant initialization ($\gamma_{0,i}\in\{0.5,1.0\}$), and (3) thresholded initialization ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,0.1)$).
-Figure 6(b) demonstrates that while all variants eliminate loss spikes, convergence quality varies within 0.36%.
-The default strategy achieves optimal final loss (2.9708 vs 2.9725 for next-best), showing that GlobalGC-guided warm-up better preserves parameter update consistency than direct initialization.
-This establishes the importance of phased initialization for gradient norm adaptation.
-Locality Granularity.
-We analyze clipping granularity across three implementation levels.
-Global-level clipping applies a single threshold for the entire model, while parameter-wise adaptation operates per individual parameter.
-On GPT-2 345M (Figure 5(a)), global clipping reduces but fails to eliminate spikes (1 event vs 0 for parameter-wise), with 0.25% (2.9639 vs 2.9712) higher final loss.
-As model scale increases to 7B parameters, TensorParallel distributed training becomes necessary for efficient optimization.
-We therefore evaluate shard-wise adaptation on Llama-2 7B, where parameters are partitioned across devices and clipping operates per-shard.
-Figure 5(b) reveals that while shard-wise clipping eliminates spikes, it introduces parameter update inconsistencies across devices, increasing loss variance by 4.3% (2.3836 vs 2.28) compared to parameter-wise granularity.
-This degradation stems from independent clipping of parameter shards disrupting holistic parameter updates.
-These results establish parameter-wise granularity as optimal, achieving precise adaptation while maintaining parameter update coherence across distributed systems.
-Figure 5: Locality granularity analysis: (a) Global vs parameter-wise clipping on GPT-2 345M model demonstrates spike elimination capability, (b) Distributed shard-wise clipping on 7B model reveals parameter integrity requirements.
-Figure 6: (a) Large rate tolerance on GPT-2 345M: AdaGC enables stable training at 10$\times$ learning rate (left), outperforms baselines (mid), and accelerates convergence through rate scaling (right).
-(b) EMA initialization strategies.
-5.4 Exploring AdaGC’s Capabilities
-We demonstrate AdaGC’s potential advantages over global gradient clipping through three critical dimensions: large learning rate tolerance, batch size scalability, and optimizer compatibility.
-Large Learning Rate Tolerance
-Figure 6(a) analyzes training stability under aggressive learning rate regimes using GPT-2 345M.
-At $3\times 10^{-3}$ learning rate (10$\times$ baseline), AdaGC maintains stable training (0 loss spikes) while achieving 2.52% lower final loss than GlobalGC (left panel).
-Comparative analysis with stabilization baselines (middle panel) reveals AdaGC’s unique capability: 1.52% lower training loss than GlobalGC+ScaledEmbed (best competitor) without spike mitigation overhead.
-The scaling law analysis (right panel) suggests a noticeable convergence acceleration with increasing learning rates (from $3\times 10^{-4}$ to $3\times 10^{-3}$).
-This indicates that learning rate scaling may contribute to a faster training process, albeit not with a perfect linear relationship.
-Batch Size Scalability
-Figure 7 validates AdaGC’s effectiveness under scaled batch training paradigms.
-With batch sizes up to 8192 (16$\times$ baseline) and proportional learning rate scaling, AdaGC maintains perfect stability (0 spikes vs 100% spike rate for GlobalGC) while reducing final perplexity by 1.28% - 19.55% across configurations.
-The Llama-2 Tiny experiments confirm cross-architecture effectiveness, showing 6.19% lower perplexity than GlobalGC at a batch size of 8192.
-This demonstrates AdaGC’s capability to enable efficient large-batch distributed training without stability compromises.
-Additional details can be found in Figure 9 and Figure 10 in Appendix D.
-(a) bs = $512$, lr = $3.0\times 10^{-4}$
-(b) bs = $2048$, lr = $6.0\times 10^{-4}$
-(c) bs = $4096$, lr = $8.5\times 10^{-4}$
-(d) bs = $8192$, lr = $1.2\times 10^{-3}$
-Figure 7: AdaGC’s large-batch scalability on GPT-2 345M: Maintains 0 loss spikes (vs GlobalGC’s 100% rate) with batch sizes up to 8192 (16$\times$ baseline).
-Optimizer Compatibility
-Figure 8 demonstrates AdaGC’s generalization across optimization frameworks.
-Integrated with Lion, AdaGC achieves comparable training loss to GlobalGC (1.9912 vs 2.0165) while improving ImageNet-1K top-1 accuracy by 0.16 percentage points (40.81% vs 40.65%).
-When combined with StableAdamW (which explicitly avoids gradient clipping), AdaGC provides additional 19.23% training loss reduction (2.0523 vs 2.5412) and 0.96 percentage point accuracy gain (39.71% vs 38.75%), establishing complementary benefits.
-These results position AdaGC as a universal stabilization component for modern optimization paradigms.
-Figure 8: AdaGC’s optimizer compatibility on CLIP ViT-Base: With Lion, achieves 0.16pp higher accuracy (40.81% vs 40.65%) at comparable loss; Combined with StableAdamW, reduces loss by 19.23% (2.0523 vs 2.5412) with 0.96pp accuracy gain.
-Demonstrates universal stabilization across optimization frameworks.
-6 Discussion
-6.1 Socio-Economic Value
-Pre-training large models entails substantial costs and is vulnerable to significant economic losses due to instabilities in the training process.
-Loss spikes, typically addressed by resuming training from the latest checkpoint as illustrated by the PaLM (Chowdhery et al., 2023) approach, can have severe economic impacts.
-According to the MegaScale (Jiang et al., 2024) report, the average recovery time from an interruption is about 15 minutes.
-Considering the Llama-3  [^1] https://ai.meta.com/blog/meta-llama-3/ model, which utilizes 16,000 NVIDIA H100 GPUs priced between \$2.29 to \$2.49 per hour  [^2] https://lambdalabs.com/, each interruption due to a Loss Spike can lead to direct economic losses ranging from approximately \$9,160 to \$9,960.
-Our proposed adaptive gradient clipping method effectively reduces the frequency of Loss Spikes, thereby enhancing training stability and minimizing economic waste.
-6.2 Limitations
-Resource constraints prevented this study from conducting long-term training experiments on the Llama-2 open-source model.
-However, our internal business validations indicate that the method remains effective throughout the mid to late stages of training.
-Future plans, resources permitting, include conducting comprehensive experiments on the Llama-2 model and sharing the findings with the academic community.
-7 Conclusion
-This paper systematically addresses the critical challenge of loss spikes in large model pre-training through gradient dynamics analysis.
-Our key insight reveals two limitations in conventional approaches: temporal gradient norm decay requires adaptive thresholds, while parameter-specific gradient heterogeneity demands localized control.
-These findings motivate AdaGC, which integrates exponential moving average-based threshold adaptation with per-parameter clipping while maintaining Adam-comparable $\mathcal{O}(1/\sqrt{T})$ convergence rate - proving local clipping preserves theoretical convergence properties.
-Comprehensive experiments demonstrate AdaGC’s dual benefits: complete loss spike elimination across LLMs (3.5% WikiText perplexity reduction on Llama-2 7B) and LVLMs (25% faster CLIP convergence than StableAdamW), coupled with enhanced training stability and final model quality.
-The consistent success across architectures and optimizers validates that jointly addressing temporal adaptation and spatial parameter characteristics provides an effective stabilization paradigm for large-scale training.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A Convergence Proof
-In this section, we provide the necessary assumptions and lemmas for the proofs of Theorem 4.1.
+Theorem F.1 shows that even with local clipped gradient, Adam with AdaGC can converge at the same rate as vanilla Adam (Kingma & Ba, 2014).
+In the following, we provide the necessary assumptions and lemmas for the proofs of Theorem F.1.
 Notations
@@
 Given a positive vector $\hat{\eta}_{t}$, it will be helpful to define the following weighted norm: $\|v_{t}\|^{2}_{\eta_{t}}=\langle v_{t},\hat{\eta}_{t}v_{t}\rangle=\sum_{k=1}^{d}\hat{\eta}_{t,k}|{v}_{t,k}|^{2}$.
-Assumption A.1.
+Assumption F.2.
 The function f is lower bounded by $\underline{f}$ with L-Lipschitz gradient.
-Assumption A.2.
+Assumption F.3.
 The gradient estimator $g$ is unbiased with bounded norm, e.g,
 $\begin{split}&\mathbb{E}[g|x_{t}]=\nabla f(x_{t}),\ \|g_{t}\|\leq G.\end{split}$
-Assumption A.3.
+Assumption F.4.
 The coefficient of clipping $h_{t,i}$ is lower bounded by some $h_{0}>0$.
-Assumption A.4.
+Assumption F.5.
 $\|g_{t}-\nabla f(x_{t})\|\leq p\|\nabla f(x_{t})\|$ holds for some $p<1$ and for all t.
-Remark A.5.
-Assumption A.1 and Assumption A.2 are widely used in the proof of optimization algorithm with adaptive learning rates (Reddi et al., 2018).
-Assumption A.3 is because the gradient norm changes slowly when training the neural network, and the last assumption holds when the batch size is large enough.
-Lemma A.6.
+Remark F.6.
+Assumption F.2 and Assumption F.3 are widely used in the proof of optimization algorithm with adaptive learning rates (Reddi et al., 2018).
+Assumption F.4 is because the gradient norm changes slowly when training the neural network, and the last assumption holds when the batch size is large enough.
+Lemma F.7.
 Let $\zeta:=\beta_{1}^{2}/{\beta_{2}}$. We have the following estimate
@@
 The proof is completed. ∎
-Lemma A.7.
+Lemma F.8.
 The following estimate holds
@@
 Then, $\|\Delta_{t}\|^{2}=\|\frac{\alpha_{t}m_{t}}{\sqrt{v_{t}}+\epsilon}\|^{2}\leq\frac{G^{2}}{\epsilon}\alpha_{t}^{2}$ by using the definition of $\Delta_{t}$.
-Therefore, $\small\sum_{t=1}^{T}\|\Delta_{t}\|^{2}\leq\frac{G^{2}}{\epsilon}\sum_{t=1}^{T}\frac{\alpha^{2}}{T}=\frac{G^{2}\alpha^{2}}{\epsilon}$.
-Lemma A.8.
-With the Assumption A.3 and A.4, it holds that
+Therefore, $\sum_{t=1}^{T}\|\Delta_{t}\|^{2}\leq\frac{G^{2}}{\epsilon}\sum_{t=1}^{T}\frac{\alpha^{2}}{T}=\frac{G^{2}\alpha^{2}}{\epsilon}$.
+Lemma F.9.
+With the Assumption F.4 and F.5, it holds that
 $\mathbb{E}\left\langle\nabla f\left(\theta_{t}\right),\hat{\eta}_{t}\hat{g}_{t}\right\rangle\geq h_{0}\mathbb{E}\left\|\nabla f\left(\theta_{t}\right)\right\|_{\hat{\eta}_{t}}^{2}.$
 Proof.
-According to Assumption A.4, it holds that
+According to Assumption F.5, it holds that
 $\begin{split}\left\langle\nabla_{i}f\left(\theta_{t}\right),g_{t,i}\right\rangle&=-\frac{1}{2}\left(\left\|\nabla_{i}f\left(\theta_{t}\right)-g_{t,i}\right\|^{2}-\left\|\nabla_{i}f\left(\theta_{t}\right)\right\|^{2}-\left\|g_{t,i}\right\|^{2}\right)\\ &\geq(1-p^{2})\left\|\nabla_{i}f\left(\theta_{t}\right)\right\|^{2}\geq 0.\end{split}$
@@
 Let $\hat{v}_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\delta_{t}^{2}$, where $\delta_{t}^{2}=\mathbb{E}_{t}\left[\hat{g}_{t}^{2}\right]$ and let $\hat{\eta}_{t}=\alpha_{t}/\sqrt{\hat{v}_{t}+\epsilon}$.
-Lemma A.9.
+Lemma F.10.
 Let $M_{t}=\mathbb{E}\left[\langle\nabla f(\theta_{t}),\Delta_{t}\rangle+L\|\Delta_{t}\|^{2}\right]$. Let $\alpha_{t}=\alpha/\sqrt{T}$ and $\beta_{2}=1-\beta/T$. Then, for $T\geq 1$ we have
@@
 For simplicity, we denote
-$\begin{split}&A_{t}^{1}=\left(1-\beta_{1}\right)\sqrt{\hat{\eta}_{t}}\hat{g}_{t}\frac{\left(1-\beta_{2}\right)\left(\sigma_{t}^{2}-\hat{g}_{t}^{2}\right)}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}}\right)}\\ &A_{t}^{2}=\beta_{1}m_{t-1}\sqrt{\hat{\eta}_{t}}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\\ &A_{t}^{3}=\beta_{1}\sqrt{\hat{\eta}_{t}}m_{t-1}\frac{(1-\beta_{2})^{2}\hat{g}_{t}^{2}(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}})}\\ &A_{t}^{4}=\beta_{1}\sqrt{\hat{\eta}_{t}}m_{t-1}\frac{(1-\sqrt{\beta_{2}})\epsilon}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}\\ &A_{t}^{5}=\beta_{1}\sqrt{\hat{\eta}_{t}}m_{t-1}\frac{(1-\sqrt{\beta_{2}})(1-\beta_{2})\epsilon(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)}\\ \end{split}$
+[…]
 Then, we obtain
@@
 For the rest of the terms, it holds that
-$\begin{split}\mathbb{E}\langle\nabla f(\theta_{t}),-\sqrt{\hat{\eta}_{t}}A_{t}^{1}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{1}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{2}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{2}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{3}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{3}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{4}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{4}\right\|^{2}\\ \mathbb{E}\langle\nabla f(\theta_{t}),+\sqrt{\hat{\eta}_{t}}A_{t}^{5}\rangle\leq\frac{h_{0}(1-\beta_{1})}{10}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}+\frac{5}{2(1-\beta_{1})h_{0}}\left\|A_{t}^{5}\right\|^{2}\\ \end{split}$
+[…]
 On the other hand, it holds that
@@
 $\begin{split}\sum_{t=1}^{T}M_{t}&\leq\sum_{t=1}^{T}\sum_{i=1}^{t}\sqrt{\zeta}^{t-i}N_{i}-\frac{(1-\beta_{1})h_{0}}{2}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\\ &\leq\frac{1}{1-\sqrt{\zeta}}\sum_{t=1}^{T}N_{t}-\frac{(1-\beta_{1})h_{0}}{2}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\\ &\leq\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}.\end{split}$
-Lemma A.10.
+Lemma F.11.
 Let $\tau$ be randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities $p_{\tau}=\frac{1}{T}$. We have the following estimate:
@@
 Thus, the desired result is obtained. ∎
-Theorem A.11.
+Theorem F.12.
 Let $\{\theta_{t}\}$ be a sequence generated by AdaGC for initial values $\theta_{1}$ and $m_{0}=v_{0}=0$.
-Assumptions A.1 to A.4 hold.
-With the hyperparameters $\alpha_{t}=\alpha/\sqrt{T}$,$\beta_{2}=1-\beta/T$ and $\zeta=\beta_{1}^{2}/\beta_{2}<1$.
+Assumptions F.2 to F.5 hold.
+With the hyperparameters $\alpha_{t}=\alpha/\sqrt{T}$$\beta_{2}=1-\beta/T$ and $\zeta=\beta_{1}^{2}/\beta_{2}<1$.
 Let $\tau$ be randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities.
@@
 $\mathbb{E}\|\nabla f(\theta_{\tau})\|^{2}\leq\frac{C}{\sqrt{T}}$
-where $C=\frac{\sqrt{G^{2}+\epsilon d}}{\alpha}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)$ and $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right)$
+where $C=\frac{\sqrt{G^{2}+\epsilon d}}{\alpha}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)$ and $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right)$.
 Proof.
@@
 $\begin{split}\mathbb{E}\left[\|\nabla f(\theta_{\tau}\|^{2}\right]&\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\right]\\ &\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\left(f(\theta_{1})-\mathbb{E}[f(\theta_{T+1})]+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)\\ \ &\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)\end{split}$
-Appendix B LLMs Hpyer-Parameters
-Table 2 shows the hyper-parameters used in our experiments on LLMs.
-Note that Llama-2 Tiny is a smaller version of the Llama-2 model, designed based on the network configuration of the GPT-2 model with 345 million parameters.
-We named it Llama-2 Tiny to facilitate our initial experimental exploration.
-Table 2: Hyper-parameters used in our LLMs experiments.
-$\lambda_{abs}$ represents the absolute global clipping threshold of GlobalGC.
-$\lambda_{rel}$ and $\beta$ represent the relative clipping threshold and the momentum of our AdaGC, respectively.
-Model LLaMA-tiny LLaMA-7B LLaMA-13B GPT-2
-Precision bfloat16 bfloat16 bfloat16 bfloat16
-Layer num 24 32 40 24
-Hidden dim size 1024 4096 5120 1024
-FFN dim size 2371 11008 13824 4096
-Attention heads 16 32 40 16
-Sequence length 2048 2048 2048 2048
-Batch size 512 2048 2048 512
-iterations 36000 9000 9000 30000
-Learning rate $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$
-Lr decay style cosine cosine cosine cosine
-Warmup iterations 2000 2000 2000 300
-Weight decay 0.1 0.1 0.1 0.1
-$\lambda_{abs}$ 1.0 1.0 1.0 1.0
-$\lambda_{rel}$ 1.05 1.05 1.05 1.05
-$\beta$ 0.98 0.98 0.98 0.98
-Appendix C Experimental Details for CLIP
-In addition to pre-training tasks for Large Language Models (LLMs), we have conducted pre-training for large-scale Vision-Language Models (LVLMs), with a specific focus on the widely acknowledged CLIP model (Radford et al., 2021).
-We train the CLIP models on the LAION-400M (Schuhmann et al., 2021) dataset and conduct zero-shot evaluations on the ImageNet (Russakovsky et al., 2015) dataset.
-We opt for the ViT-Base configuration, which contains 151 million parameters.
-The ViT-Base model is trained over 20K steps, observing 320M samples.
-In alignment with (Wortsman et al., 2023a), we implement patch-dropout 0.5 (Li et al., 2023), learning rate 0.002 and weight decay 0.2, with the initial 5K steps dedicated to linear warmup (Goyal et al., 2017) and the subsequent steps follow a cosine decay pattern (Loshchilov & Hutter, 2016).
-Appendix D Results of Batch Size Scalability Experiments
-(a) batch size = $512$, lr = $3.0\times 10^{-4}$
-(b) batch size = $2048$, lr = $6.0\times 10^{-4}$
-(c) batch size = $4096$, lr = $8.5\times 10^{-4}$
-(d) batch size = $8192$, lr = $1.2\times 10^{-3}$
-Figure 9: AdaGC’s large-batch scalability on GPT-2 345M: Maintains 0 loss spikes (vs GlobalGC’s 100% rate) with batch sizes up to 8192 (16$\times$ baseline) and proportional learning rate scaling, achieving 1.28%-19.55% perplexity reduction.
-(a) batch size = $512$, lr = $3.0\times 10^{-4}$
-(b) batch size = $2048$, lr = $6.0\times 10^{-4}$
-(c) batch size = $4096$, lr = $8.5\times 10^{-4}$
-(d) batch size = $8192$, lr = $1.2\times 10^{-3}$
-Figure 10: AdaGC’s large-batch scalability on Llama-2 Tiny: Cross-architecture validation shows 6.19% lower perplexity at 8192 batch size.